### v1

# ID-Guard: A Universal Framework for Combating Facial Manipulation via Breaking Identification

Abstract
The misuse of deep learning-based facial manipulation poses a potential threat to civil rights. To prevent this fraud at its source, proactive defense technology was proposed to disrupt the manipulation process by adding invisible adversarial perturbations into images, making the forged output unconvincing to the observer. However, their non-directional disruption of the output may result in the retention of identity information of the person in the image, leading to stigmatization of the individual. In this paper, we propose a novel universal framework for combating facial manipulation, called ID-Guard. Specifically, this framework requires only a single forward pass of an encoder-decoder network to generate a cross-model universal adversarial perturbation corresponding to a specific facial image. To ensure anonymity in manipulated facial images, a novel Identity Destruction Module (IDM) is introduced to destroy the identifiable information in forged faces targetedly. Additionally, we optimize the perturbations produced by considering the disruption towards different facial manipulations as a multi-task learning problem and design a dynamic weights strategy to improve cross-model performance. The proposed framework reports impressive results in defending against multiple widely used facial manipulations, effectively distorting the identifiable regions in the manipulated facial images. In addition, our experiments reveal the ID-Guard’s ability to enable disrupted images to avoid face inpaintings and open-source image recognition systems.
Index Terms:
Deepfake, facial manipulation, adversarial attack, identity protection, multi-task learning.
I Introduction
The spread of false information in communities has been a longstanding concern, presenting a potential threat to civil rights and social security. In recent years, the advancement and deployment of generative deep neural networks (DNNs) have exacerbated this issue, with facial manipulation serving as a notable example. This technology enables end-to-end manipulation of facial attributes or identity of images/videos. Malicious actors, for instance, exploit forged images to generate and circulate misleading news [1, 2] or perpetrate online fraud [3]. Although the re-training of these methods is proved challenging due to high computing power requirements and technical thresholds, users can easily download accessible pre-trained models from open-source platforms such as Github  [^1] https://github.com, Hugging Face  [^2] https://huggingface.com, and TensorFlow Hub  [^3] https://www.tensorflow.org to implement the forgery [4]. This greatly reduces the cost of creating fake examples, thereby expediting the proliferation of false information on social media. Therefore, an urgent need exists to develop proactive and efficient defense mechanisms.
Figure 1: Illustration of the impact of malicious propagation of facial manipulation samples. Fakes will lead to rumor spreading and insufficient distortion of faces by traditional defense methods will cause face stigmatization. Our method disrupts the observer’s identification of the identity in the sample and thus adequately protects the individual’s rights.
To mitigate the aforementioned threats, substantial research efforts have been directed toward proactive defense mechanisms against facial manipulation in recent times. Unlike passive detection methods [5, 6, 7, 8, 9], proactive defense algorithms [10, 11, 12, 4, 13, 14, 15, 16, 17, 18, 19] are crafted to thwart fraudulent activities at their source. These algorithms induce visual distortions in facial manipulation outputs by introducing imperceptible adversarial perturbations [20] into face images. However, the disruption caused by existing algorithms to manipulated images is often non-directional, leading to the preservation of individuals’ identifiable information in the disrupted images. When stubborn malicious users insist on uploading these incompletely disrupted forged images to social media, concerns such as face stigmatization may arise [16]. As illustrated in Fig. 1, applying the traditional proactive defense algorithm mentioned above to safeguard photographs of the renowned movie star Jackie Chan still allows ordinary Internet users to discern his identity readily. Insufficient distortion of facial images makes the disrupted output a “spoof”, leaving the victim’s reputation and rights inadequately protected. Notably, as shown in Fig. 2, above insufficiently disrupted facial examples also face two potential threats: 1) The residual identifiable information in the samples makes them still likely to be recognized by commercial facial recognition systems. The stigmatization problem is aggravated by the fact that some entertainment applications automatically identify and push images of celebrities; 2) Some malicious forgers with rich technical capabilities may restore the fake images that are not seriously distorted to continue to commit fraud.
Figure 2: Illustration of potential threats to the insufficiently disrupted facial example. Challenges come primarily from commercial face recognition systems and face inpainting algorithms.
To address the above concerns, in this paper, we propose a proactive defense framework called ID-guard. This framework requires only a single forward pass of an image reconstruction network to produce the universal adversarial perturbation to combat multiple open-source facial manipulation algorithms, as shown in Fig. 3. To destroy the identifiable semantic information in forged images and prevent spoofing of individuals, a novel Identity Destruction Module (IDM) is introduced. IDM enables the generated protective perturbations to focus on disrupting key regions of the individual’s face in an image to prevent an observer from recognizing his/her identity.
In addition, the universality of the adversarial perturbations is crucial in practical application scenarios [12] due to the facial manipulation methods used by the forger are unknown and uncontrollable. To improve the cross-model universality of perturbations, a dynamic weighting strategy is proposed during the training of the perturbation generator. Specifically, the robustness of different facial manipulations against attacks varies due to differences in model structure and complexity. Simply treating the adversarial loss of attacking different models as equally weighted when training the perturbation generator will result in the produced perturbations being biased towards facial manipulations that are easy to attack. Therefore, we consider the attacks against different models as a multi-task learning problem, which enables the loss weights to be dynamically adjusted during the training process to achieve well-balanced cross-model performance. A gradient prior perturbation strategy is also introduced to guide the perturbation generator to improve the stability of training and accelerate convergence.
As expected and will be verified by experiments, the proposed ID-guard can effectively distort recognizable regions in facial images forged by various open-source manipulation algorithms, thereby preventing observers and face recognition systems from recognizing their identities and circumventing face inpaintings. In summary, the contributions of our work are summarized as follows:
  1. 1.
    We propose a novel general adversarial perturbation generation framework to prevent facial manipulation from stigmatizing individuals, called ID-guard. Only a single forward pass of the generator is required to quickly produce perturbations that can disrupt various facial manipulations.
  2. 2.
    To fully disrupt the manipulation results so that the identities of individuals in the images cannot be identified, an Identity Destruction Module is proposed to guide the generated perturbations to focus on distorting identity-related semantic information. This module can further alleviate concerns about commercial face recognition systems and image inpainting algorithms.
  3. 3.
    To improve the universality of the generated adversarial perturbations, we implement an attack on multiple facial manipulations by solving a multi-task learning problem and designing a dynamic weighting strategy. To improve the stability of the generator, we introduce a gradient prior perturbation strategy.
The remainder of this paper is organized as follows: Related works on Deepfakes and proactive defense are reviewed in Section II. Section III presents the details of our method. The experimental results and analysis are provided in Section IV, followed by the conclusion in Section V.
Figure 3: The accessible pre-trained models can be easily downloaded from open-source platforms to implement forgeries. For a given image, the proposed ID-guard can generate universal perturbations for defense against multiple open-source facial manipulations through a single forward propagation of an image reconstruction network.
II Related Works
II-A Facial Manipulation
Facial manipulation aims at constrained modification of the face of a given image/video to present the desired visual content, such as identity, expression, age, hair color, etc. of the character. In recent years, thanks to the great success of Generative Adversarial Networks (GANs) in image synthesis, many GAN-based algorithms [21, 22, 23, 24, 25] with different designs and constraints have been proposed to solve facial manipulation [26]. Some contributors choose to open source their work on the public platform, including pre-training models and running scripts, which greatly lowers the threshold for users to produce high-quality and high-fidelity fake images/videos.
II-B Facial Manipulation Disruption
Some recent works have achieved proactive defense against facial manipulation by injecting adversarial perturbations into the image. Ruiz et al.[10] and Yeh et al.[11] achieved disruption of the facial manipulation by deriving gradient-based adversarial perturbations on the target model. Works including [18] and [19] have significantly improved the robustness of adversarial perturbations for protecting personal images. However, the common drawbacks of these methods are higher computational overhead and lower work efficiency, due to they involve solving the model-specific perturbation for each image. Methods represented by [12, 4, 17] and [15] have studied cross-model universal perturbations, which to some extent alleviate overhead issues and practicality limitations. However, they ignore the fact that different facial manipulation models have differences in robustness and gradient direction, and thus the performance of the produced generic perturbations is uneven across models. Furthermore, as discussed above, these algorithms do not consider the problem of face stigmatization due to unconstrained disruptions. Zhai et al. [16] proposed a method to embed specific warning patterns on the generated fake images to solve this problem. Unlike them, the proposed ID-guard directly distorts the facial recognition area of fake images.
II-C Multi-task Learning
One of the effective routes to achieve multi-task learning is to dynamically weight the losses of different tasks according to their learning stages or the difficulty of learning. Sener et al. [27] pointed out that multi-task learning can be regarded as a multi-objective optimization problem, aiming to find the Pareto optimal solution to optimize the performance of multiple tasks. A representative method that has been proven effective and widely used is the multiple gradient descent algorithm (MGDA) [28]. Some heuristic works [29, 30, 31, 32, 33] measured the difficulty of a task based on the order of magnitude or change rate of the loss value, and then dynamically adjusted the weights of different tasks to obtain balanced performance. In this work, we further explore the potential of integrating multi-task learning strategies into across-model universal perturbation generation.
III Methodology
In this section, the specific design and implementation details of the proposed ID-guard framework are elaborated. For clarity, we first introduce the overview of the framework and the definition of notations.
III-A Preliminaries
III-A1 Adversarial Attacks against Facial Manipulations
Given a natural image $x\in\mathbb{R}^{3\times H\times W}$, the pre-trained facial manipulation model $\mathcal{M}$ translates it into a forgery $y=\mathcal{M}(x)$. The goal of the defender is to find a small adversarial perturbation $\delta$ that makes the facial manipulation model fail to manipulate the perturbed adversarial image $x_{adv}=x+\delta$. The solution to the $\delta$ can be formulated as a maximization optimization problem as follows
$\begin{split}&\max_{\eta}\mathcal{D}(\mathcal{M}(x),\mathcal{M}(x+\delta)),\\ &s.t.\quad\|\delta\|_{\infty}\leq\epsilon,\end{split}$ (1)
where $\mathcal{D}$ is the metric of the distance between images and its design is our focus and will be introduced in subsequent sections. $\epsilon$ is the infinite norm bound used to restrict the perturbation. The existing mainstream adversarial perturbation derivation patterns can be roughly divided into two categories: gradient-based methods [34, 35] and generation-based methods[36, 37]. Compared to the former, the latter is more flexible and faster, so we choose the generation-based approach to drive the perturbation derivation of the proposed ID-guard. Specifically, we train a perturbation generator $\mathcal{G}$ with the following optimization objective to produce the desired adversarial perturbation:
$\begin{split}&\max_{{\theta}_{\mathcal{G}}}\mathbb{E}(\mathcal{D}(\mathcal{M}(x),\mathcal{M}(x+\mathcal{G}(x)))),\\ &s.t.\quad\|\mathcal{G}(x)\|_{\infty}\leq\epsilon\end{split}$ (2)
$\mathcal{G}$ is designed as a Resnet architecture [38], and the advantages of this design will be discussed in subsequent experiments.
III-A2 Cross-Model Universal Perturbations
We hope that the adversarial perturbation generated for a certain image can be effective for a set of open-source pre-trained facial manipulation models $\mathcal{S}_{\mathcal{M}}=\left\{\mathcal{M}_{1},\mathcal{M}_{2},\dots,\mathcal{M}_{N}\right\}$. Hence, the optimization objective in Eq. (2) can be rewritten as:
$\begin{split}&\max_{{\theta}_{\mathcal{G}}}\sum_{k=1}^{N}\lambda_{k}\mathbb{E}(\mathcal{D}(\mathcal{M}_{k}(x),\mathcal{M}_{k}(x+\mathcal{G}(x)))),\\ &s.t.\quad\|\mathcal{G}(x)\|_{\infty}\leq\epsilon\end{split}$ (3)
where $\mathcal{S}_{\lambda}=\left\{\lambda_{1},\lambda_{2},\dots,\lambda_{N}\right\}\in\mathbb{R}$ is a set of weights got based on the robustness of each target model. Solving the effective $\mathcal{S}_{\lambda}$ is one of the key tasks of this work so that the generated adversarial perturbations have a balanced defense performance against different facial manipulations. This part will be introduced in detail in Section III-C.
Figure 4: Illustration of the proposed ID-guard framework. The perturbation generator takes a natural image $x$ as input and requires only one forward propagation to generate a cross-model adversarial perturbation dedicated to the input face that can be used to defend against multiple facial manipulations. In the training phase, ID-guard consists of three modules, including the identity destruction module, the dynamic weighting strategy, and the gradient prior perturbation strategy. The notation descriptions are shown in the upper right corner for reference.
III-B Identity Destruction Module
As mentioned above, the design of the distance metric $\mathcal{D}$ is very critical. Traditional methods [10, 11, 39, 12] generally choose Mean Squared Error (MSE) as a proxy for distance measurement. However, this approach causes non-directional distortion of forged images and fails to ensure that the identity of the distorted face cannot be recognized. As an improvement, we introduce the Identity Destruction Module (IDM). As shown in Fig. 4, the IDM consists of three sub-modules, which will be introduced separately next.
III-B1 Mask Loss
First, we consider using face masks to limit the regions of image distortion by adversarial perturbations. The designed mask is two-fold: 1) The binary mask is used to restrict image distortion to areas of facial components including the eyes, nose, mouth, and eyebrows, which are proven to play an important role in identity recognition by the human eye [40, 41, 42]; 2) The heatmap mask weights the face distortion loss at the pixel level, making the perturbation pay more attention to the important feature areas of the face. In this work, the heatmap of each image is obtained by solving Grad-SAM [43] on VGGFace [44]. This design will also facilitate distorted images against commercial facial recognition systems. Hence, for the facial manipulation model $\mathcal{M}_{k}$, the mask loss can be formulated as:
$\mathcal{L}_{k}^{mask\_bin}=\|\mathcal{M}_{k}(x)\odot m^{bin}-\mathcal{M}_{k}(x+\mathcal{G}(x))\odot m^{bin}\|_{2},$ (4)
$\mathcal{L}_{k}^{mask\_hm}=\|\mathcal{M}_{k}(x)\odot m^{hm}-\mathcal{M}_{k}(x+\mathcal{G}(x))\odot m^{hm}\|_{2}$ (5)
where $m^{bin}$ and $m^{hm}$ denote the binary mask and heatmap mask of the natural image $x$, respectively. Note that these masks are only computed during the perturbation generator training stage to constrain the distortion region and are not used in the inference process. $\odot$ indicates the element-wise multiplication.
III-B2 Identity Loss
In addition to pixel-level constraints, further consideration is given to maximizing the identity similarity of the manipulation results on the original and adversarial images. We achieve this by maximizing the cosine distance between their identity embeddings:
$\mathcal{L}_{k}^{ID}=\cos(ID(\mathcal{M}_{k}(x)),ID(\mathcal{M}_{k}(x+\mathcal{G}(x))))$ (6)
where $ID(\cdot)$ denotes the Arcface net [45] designed to extract high-quality features from facial images and embed them into a low-dimensional space where the distance between different embeddings corresponds to the similarity between faces.
III-B3 Feature Loss
Learning from [17] that the process of attribute or identity embedding for end-to-end facial manipulation varies, but the feature extraction process is similar. In addition, feature-level perturbations can retain their effective components in network transmission to a greater extent [16]. Therefore, a feature loss that enables the generator to focus on destroying feature-level faces is incorporated into the IDM to improve the effectiveness and transferability of produced perturbations:
$\mathcal{L}_{k}^{feat}=\|E_{k}(x)\odot LI(m^{bin})-E_{k}(x+\mathcal{G}(x))\odot LI(m^{bin})\|_{2}$ (7)
where $E_{k}$ is the feature extraction module of $\mathcal{M}_{k}$, which is defined here as the upsampling network of each model. $LI(\cdot)$ represents the linear interpolation operation to make the binary mask and the extracted feature map consistent in image size.
In summary, given a pre-trained facial manipulation model $\mathcal{M}_{k}\in\mathcal{S}_{\mathcal{M}}$, the attacking loss against it can be formulated as:
$\mathcal{L}_{k}=\mathcal{L}_{k}^{mask\_bin}+\mathcal{L}_{k}^{mask\_hm}+\mathcal{L}_{k}^{ID}+\mathcal{L}_{k}^{feat}$ (8)
The loss function for training the perturbation generator $\mathcal{G}$ is a linear combination of attacking losses of facial manipulation models:
$\mathcal{L}=\lambda_{1}\cdot\mathcal{L}_{1}+\lambda_{2}\cdot\mathcal{L}_{2}+\dots+\lambda_{N}\cdot\mathcal{L}_{N}$ (9)
The definition of $\lambda_{k}$ is shown in Eq. (3), and the use of the proposed dynamic weight strategy to solve it will be introduced next.
III-C Dynamic Weight Strategy
We regard the attack against different facial manipulation models in Eq. (9) as a multi-task learning problem to solve the weight set $\mathcal{S}_{\lambda}$. As shown in Fig 4, the proposed dynamic weighting strategy is two-fold. First, a classical but effective multi-task learning strategy, i.e., MGDA [28], is integrated into the proposed ID-guard framework in order to dynamically adjust the weights to optimize the gradients of multiple tasks during the generator training phase. Specifically, we follow [27]  [^4] https://github.com/isl-org/MultiObjectiveOptimization to solve this multi-objective optimization problem and use it as a baseline dynamic weighting strategy (noted as S-I).
On the other hand, we introduce a strategy that combines prior weights and dynamic weights (noted as S-II). For the $\lambda_{k}$ of the attack loss in Eq. (9), refine it to
$\lambda_{k}^{t}=\alpha_{k}\times\beta_{k}^{t}$ (10)
where $t$ indexes the iteration steps, $\alpha_{k}=10^{n}\,(n\in\mathbb{Z})$ is a prior order of magnitude weight for the attack loss, which is determined based on the adversarial robustness of the corresponding facial manipulation model. Here, a heuristic is used to quantify the adversarial robustness: on a mini-batch of images, the equivalent setting PGD [35] is used to derive a gradient-based adversarial attack against facial manipulation models, and the $L_{2}$ distance of the attack is calculated. $\beta_{k}^{t}$ is dynamically adjusted during the training stage. Inspired by [31], we would like to make the harder-to-learn task, i.e., the harder-to-attack facial manipulation model, have higher weights during training. Hence, $\beta_{k}^{t}$ is solved as:
$\beta_{k}^{t}=-(1-\mathcal{K}_{k}^{t})\log\mathcal{K}_{k}^{t}$ (11)
where $\mathcal{K}_{k}$ represents the KPI for $\mathcal{M}_{k}$, which is an indicator of attack performance. KPI is inversely proportional to the training difficulty of the task, that is, the higher the KPI, the easier the task is to learn. Here we choose $L_{2}$ distance as the proxy of KPI. Intuitively, tasks with high KPIs are easier to learn and therefore become less weighted; conversely, tasks that are difficult to learn become more weighted. This combined design improves the stability of dynamic weights to produce adversarial perturbations with more balanced performance.
III-D Gradient Prior Perturbation
One obstacle to training adversarial perturbation generators is their lack of initial awareness of structural perturbation information. Therefore, we introduce a gradient prior perturbation strategy. Motivated by [4], we consider jointly optimizing for a global prior perturbation $\delta_{p}\in\mathbb{R}^{3\times H\times W}$ and the generator $\mathcal{G}$. Specifically, we first train a surrogate model $\mathcal{M}_{s}$ with face reconstruction capabilities, treating it as an approximate task of facial manipulation [14, 46]. Next, we use PGD [35] to derive gradient-based adversarial perturbations against $\mathcal{M}_{s}$ on a batch of face images, and average these perturbations to obtain $\delta_{p}$. More details will be introduced in IV-A5. Therefore, the overall optimization objective in Eq. (3) can be rewritten as:
$\begin{split}\max_{{\theta}_{\mathcal{G}},\delta_{p}}\sum_{k=1}^{N}\lambda_{k}\mathbb{E}&(\mathcal{D}(\mathcal{M}_{k}(x),\mathcal{M}_{k}(x+\mathcal{G}(\mathrm{cat}(x,\delta_{p}))))+\\ &\mathcal{D}(\mathcal{M}_{k}(x),\mathcal{M}_{k}(x+\delta_{p})),\\ &s.t.\quad\|\mathcal{G}(x)\|_{\infty}\leq\epsilon,\quad\|\delta_{p}\|_{\infty}\leq\epsilon\end{split}$ (12)
where $\mathrm{cat}(\cdot)$ denotes channel-wise concatenation, i.e., $\mathrm{cat}(x,\delta_{p})\in\mathbb{R}^{6\times H\times W}$. The intuition behind this design is that the gradient prior perturbation can provide the generator with rich prior gradient and structural perturbation information, thereby promoting more stable training and more efficient perturbation generation.
IV Experiments
IV-A Experimental Setup
IV-A1 Datasets
In our experiments, the CelebAMask-HQ [47] dataset is selected for training the perturbation generator. It consists of more than 30,000 face images, where each image carries semantic masks for 19 facial component categories. These fine-grained mask labels can provide support for computing the binary mask loss during the training stage. To adequately evaluate the performance as well as the generalizability of our method and the competing algorithms, we test them on three datasets including CelebAMask-HQ [47], LFW [48] and FFHQ [49].
IV-A2 Target Models
We choose five facial manipulation models including StarGAN [21], AGGAN [22], FPGAN [23], RelGAN [24] and HiSD [25] as target models to implement the attack, and they are all trained on the CelebA [50] dataset. In the experiment, for StarGAN, AGGAN, FPGAN, and RelGAN, we select black hair, blond hair, brown hair, gender, and age as editing attributes; for HiSD, five images with black hair, blond hair, brown hair, glasses, and bangs are chosen as attribute references, respectively.
IV-A3 Baselines
To demonstrate the superiority of the proposed method in face identity protection and cross-model universal performance, four advanced proactive defense methods including Disrupting [10], PG [39], CMUA [12] and IAP [15] are selected as competing algorithms. Disrupting [10] disrupts facial manipulation by iteratively solving gradient-based adversarial perturbations on the target model. PG [39] achieves adversarial perturbation generation in gray-box scenarios by attacking a surrogate model. CMUA [12] is a baseline of universal defense against multiple models. IAP [15] designs an information-containing adversarial perturbation, but we only implement its proactive distortion without involving the embedding and extraction of information to provide a fair comparison.
IV-A4 Metrics
Unlike traditional evaluation methods that calculate the $L_{2}$ distance of the whole image or the forged area between the forged and distorted outputs, we focus on measuring the difference in the facial area of the output. Specifically, we introduce $L_{2}^{face}$, which can better reflect whether the defense successfully destroys the identity information of the face image, making it unrecognizable. $L_{2}^{face}$ can be expressed as:
$L_{2}^{face}(y,\hat{y})=\frac{\sum_{i}\sum_{j}Face_{i,j}\odot(y_{i,j}-\hat{y}_{i,j})^{2}}{\sum_{i}\sum_{j}Face_{i,j}}$ (13)
where $(i,j)$ is the coordinate of pixels and $Face_{i,j}$ is the binary facial mask of the image. The pixel value of its face area is 1, otherwise it is 0. The binary facial mask is calculated by Dlib  [^5] https://pypi.org/project/dlib. Additionally, we evaluate the identity similarity (noted as ID sim. in Tables) computed by Arcface [45] between the forged and distorted outputs. Defense success rates are also considered. Previous works [10, 12] have generally determined the success of a defense by whether the $L_{2}$ distance is greater than 0.05, but this is incomplete in the task of preventing face stigmatization. As shown in Fig. 5, the distorted output in Fig. 5 (c) reports a successful defense at the $L_{2}^{face}$ distance, but it appears that it merely “blackens” the face without destroying the individual’s identity. Therefore, we propose that both $L_{2}^{face}$ distance greater than 0.05 and identity similarity less than 0.4 be satisfied to indicate successful defense, which is a more challenging evaluation. The contrast between (d) and (e) in Fig. 5 shows the necessity of considering both restrictions simultaneously.
Figure 5: Illustration of indicators that determine the success of a defense. (a) is a natural image, (b) is a forged image, (c), (d), and (e) are distorted images under different defense situations, respectively. It can be seen that when only one of the metrics, $L_{2}^{face}$ distance or identity similarity, satisfies the set conditions, it is not sufficient to break the identity of the individual in the image.
IV-A5 Implementation Details
All images used in experiments are resized to a resolution of $256\times 256$ and the pixel value is normalized to $\left[-1,1\right]$. For fairness, the bound $\epsilon$ of all competitive algorithms is restricted to 0.05 to ensure the invisibility of the perturbation. For StarGAN, AGGAN, FPGAN, RelGAN, and HiSD, we set the prior weight $\alpha$ to $\left[1,1,1,10,100\right]$, respectively, as determined by a simple pre-experiment on the gradient-based adversarial attack against them. We derive the gradient prior perturbation on 2,000 randomly selected face images from CelebAMask-HQ[47], running the PGD [35] for 10 iterations with a 0.01 step size. The perturbation generator is trained using the Adam [51] with a learning rate of 0.0001, and the batch size is 32.
TABLE I: Quantitative comparison for disrupting different target models. For each column within the same dataset. The best result is marked in bold, while the sub-optimal result is marked with an underline.
  Datasets Methods StarGAN[21] AGGAN[22] FPGAN[23] RelGAN[24] HiSD[25]
[…]
CelebA-HQ Disrupting[10] 1.047 0.023 1.000 0.114 0.479 0.292 0.134 0.369 0.472 0.021 0.753 0.001 0.004 0.839 0.001
PG[39] 0.101 0.302 0.646 0.032 0.658 0.014 0.069 0.352 0.458 0.007 0.836 0.005 0.016 0.617 0.095
CMUA[12] 0.586 0.368 0.584 0.062 0.646 0.016 0.052 0.486 0.126 0.296 0.630 0.070 0.055 0.603 0.165
IAP[4] 0.450 0.118 0.994 0.054 0.300 0.398 0.321 0.181 0.928 0.109 0.545 0.165 0.056 0.193 0.590
Ours (S-I) 0.362 0.055 1.000 0.376 0.062 1.000 0.558 0.004 1.000 0.285 0.065 0.998 0.205 0.018 1.000
Ours (S-II) 0.587 0.018 1.000 0.300 0.089 1.000 0.631 0.000 1.000 0.408 0.010 1.000 0.202 0.042 0.998
  LFW Disrupting[10] 0.956 0.062 1.000 0.142 0.443 0.412 0.126 0.380 0.546 0.023 0.075 0.011 0.004 0.849 0.000
PG[39] 0.134 0.306 0.728 0.053 0.656 0.044 0.069 0.415 0.356 0.008 0.838 0.002 0.020 0.651 0.035
CMUA[12] 0.513 0.247 0.788 0.092 0.462 0.294 0.054 0.745 0.108 0.231 0.618 0.078 0.063 0.600 0.150
IAP[4] 0.413 0.411 0.956 0.085 0.411 0.424 0.310 0.161 0.946 0.079 0.533 0.211 0.071 0.238 0.545
Ours (S-I) 0.328 0.021 1.000 0.446 0.033 1.000 0.555 0.027 0.994 0.271 0.084 0.992 0.178 0.053 0.998
Ours (S-II) 0.522 0.047 1.000 0.414 0.078 1.000 0.644 0.017 1.000 0.430 0.022 1.000 0.221 0.081 0.980
  FFHQ Disrupting[10] 0.956 0.033 1.000 0.142 0.487 0.312 0.126 0.409 0.426 0.023 0.747 0.013 0.005 0.878 0.002
PG[39] 0.134 0.328 0.628 0.053 0.708 0.002 0.069 0.424 0.360 0.008 0.839 0.000 0.020 0.683 0.108
CMUA[12] 0.515 0.346 0.624 0.092 0.635 0.028 0.054 0.512 0.168 0.231 0.666 0.043 0.063 0.656 0.120
IAP[4] 0.413 0.167 0.976 0.085 0.365 0.476 0.310 0.207 0.898 0.079 0.568 0.123 0.071 0.252 0.525
Ours (S-I) 0.328 0.053 1.000 0.446 0.089 0.994 0.557 0.019 1.000 0.271 0.112 0.973 0.178 0.050 0.965
Ours (S-II) 0.536 0.002 1.000 0.329 0.107 0.970 0.502 0.025 1.000 0.345 0.030 0.995 0.183 0.081 0.964
IV-B Comparison with Baselines
Table I summarizes the quantitative comparison of the proposed ID-guard with competitive algorithms. Our method is reported separately under two strategies, as presented in III-C. The Disrupting [10] produces gradient-based adversarial perturbations against StarGAN [21], which makes it overfit in disrupting this model at the cost of cross-model performance. The perturbation optimization of CMUA [12] and PG [39] is unconstrained and thus has limited destruction to the identity semantics. As shown in Fig. 6, the person’s identity in a distorted image under the CMUA can be still recognized. Attributed to the feature correlation measurement loss employed, IAP [15] improves the performance to destroy identification to some limit extent. In comparison, our method significantly destroys the identity semantics of images and effectively prevents the stigmatization of faces. Furthermore, the baseline methods equally weight the attack loss of different facial manipulations, leading to perturbations biased towards vulnerable models. Due to the introduction of the dynamic parameter strategy, we achieve balanced performance against various facial manipulations. For the most robust model against attack, HiSD [25], ID-guard improves the defense success rate by 50% compared with the state-of-the-art method.
On the three selected datasets, ID-guard under both strategies demonstrated superior performance. For strategy I, the MGDA algorithm is employed to adjust the weights, which has the advantage of completely eliminating the need for human intervention and prior knowledge during the training process. However, this non-intervention makes ID-guard based on strategy I exhibit an extreme effect to some extent: it performs better on the more vulnerable model (e.g., AGGAN [22]) and the more robust model (e.g., HiSD [25]), but does not provide significant improvement on models in the middle (e.g., RelGAN [24]). Moreover, additional backpropagation computation is required at each iteration when implementing MGDA, which increases the training overhead. In contrast, Strategy II-based ID-guard only additionally computes a set of KPI values, which has a minimal impact on training overhead. Strategy II maintains a set of prior parameters for balancing each model, thereby further balancing performance. The disadvantage of it is the extra work of determining the prior parameter set.
Figure 6: Visual examples of disruption to different facial manipulations. Compared with CMUA [12], which still retains a large amount of identity semantics in the distorted image, our method makes the face in the image completely unrecognizable.
IV-C Ablation Study
IV-C1 Identity Destruction Module
The identity destruction module aims to destroy the identity semantics of a face so that it cannot be correctly recognized. We delve into the impact of the three designed losses on the destruction effect. Table II and Fig. 7 present the quantitative and visual ablation results, respectively. Specifically, the three sub-modules focus on different issues. The mask loss uses two facial masks as strong constraints for the attack, thus providing a huge improvement in significantly distorting facial regions. Identity loss is a feature-level constraint that perturbs the key areas of identity recognition from a global perspective of the image. This design is important in destroying machine identification and will be introduced in detail in Section IV-D1. As shown in Fig. 7, the mask loss concentrates the distortion on the face region of the image, while the identity loss destroys the global texture. Feature loss brings overall gain, which benefits from the similarity in feature extraction of the face manipulation model. It is worth noting that the three types of losses reinforce each other to some extent.
TABLE II: Ablation results for component modules. The best result in each column is marked in bold.
[…]
#1 w/o all 0.172 0.509 0.358
#2 w/o Feature loss 0.396 0.050 0.973
#3 w/o Id. loss 0.431 0.148 0.937
#4 w/o Mask loss 0.189 0.039 0.866
#5 w/ all 0.425 0.031 0.999
Figure 7: Visual examples of ablation study of identity destruction module. Among them (a) is a natural image, (b) is a fake image, and (c)-(f) correspond to conditions #1, #3, #4, and #5 in Table II respectively.
IV-C2 Dynamic Weight Strategy
The dynamic weight strategy focuses on balancing the attack losses for different facial manipulations. We selected equivalent weight, prior weight, hard model mining (HMM) [13], and KPI as the baseline of the weight setting methods. The equivalent weight setting will cause the generated perturbations to overfit on the most vulnerable model architecture (e.g., StarGAN and FPGAN). Although HMM balances each model to a certain extent, it ignores the difference in model gradients and thus causes the degradation of average performance. Separate prior weight setting or KPI are unstable and difficult to set, so we cleverly blend the two in Strategy II and get stable training. The benefits of this are two-fold: 1) It reduces the difficulty of a prior setting, and only needs to determine a series of orders of magnitude to allow automatic optimization of parameters; 2) It makes the KPI strategy more stable. Strategy I also achieves excellent results, but the additional backpropagation makes its training more expensive.
TABLE III: Comparison of defense success rates under different optimization strategies. The best result in each column is marked in bold, while the sub-optimal result is marked with an underline.
  Optimizations StarGAN AGGAN FPGAN RelGAN HiSD Average
Equivalent weight 1.000 0.458 1.000 0.095 0.420 0.595
Prior weight 0.967 0.986 1.000 0.913 0.875 0.948
HMM 0.990 0.982 0.985 0.986 0.681 0.925
DTP 1.000 0.894 1.000 0.802 0.885 0.916
Ours (S-I) 1.000 1.000 1.000 0.998 1.000 0.999
Ours (S-II) 1.000 1.000 1.000 1.000 0.998 0.999
IV-C3 Gradient Prior Perturbation
Gradient prior perturbation aims to provide the generator with noise-like prior knowledge, thus accelerating its convergence. For comparison, the variation of training loss and defense performance at different scales with gradient a prior perturbation, with a prior random noise, and without prior knowledge are shown in Fig. 8. Both gradient prior perturbation and random noise promote the convergence of the generator, which is due to the introduction of global noise structure [4]. In terms of generator performance, methods based on gradient prior perturbation at different scales have shown the most significant defense effect, with an average improvement of 31.2% compared to random noise methods. We believe that the reason behind this is that gradient prior perturbation involves rich adversarial structural information.
(a) Changes of training loss
(b) Defense results at different scales
Figure 8: The changes of training loss and defense performance at different scales with gradient a gradient prior perturbation, with random noise, and without any prior knowledge.
IV-C4 Architecture of the Perturbation Generator
We explore the impact of different generator architectures on performance. Three mainstream architectures including Unet [52], Resnet [38] and Transformer [53] were selected as the generators of the proposed ID-guard. Fig. IV reports the defense performance of the generators for these three architectures. Compared with Unet, Resnet and Transformer architectures have achieved significant advantages. As shown in Table IV, Transformer achieved optimal performance at the expense of model parameter size, while Resnet achieved very close performance with less than 5% of its parameter size. We propose to use Resnet as the architecture for the generator of the proposed ID-guard, and the intuition behind this is that the generated perturbation can be regarded as a residual of the image.
(a) $L_{2}^{face}$
(b) Defense success rate
Figure 9: Performance comparison of perturbation generators based on different architectures.
TABLE IV: Comparison of the number of model parameters for generators based on different architectures.
  Generator architectures Number of parameters
Unet-based Generator 54,414,595
Resnet-based Generator 7,850,819
Transformer-based Generator 179,348,843
IV-D Other Evaluation
IV-D1 Misleading Facial Recognition Systems
Some social applications recognize photos uploaded by users and then add corresponding tags and use them in content recommendation systems. This can exacerbate the spread of distorted faces. Therefore, the threat of stigmatization of distorted images comes not only from the human eye but also from commercial facial recognition systems. Therefore, the threat of stigmatization of distorted images comes not only from the human eye but also from commercial facial recognition systems. As shown in Fig. 10, we evaluate the misdirection success rates of the destroyed outputs of ID-guard and competing algorithms on three mainstream face recognition systems. As can be seen, our method reports optimal results, achieving over 95% misdirection success rate on Google  [^6] https://www.google.com and StarByFace  [^7] https://starbyface.com. Baidu  [^8] https://www.baidu.com has the most robust recognition system, with CMUA [12] and PG [39] can hardly fool it, but ID-guard still causes it to recognize more than 75% of images incorrectly. The good performance of ID-guard is due to the identity loss introduced to destroy the identity recognition baseline model, which is widely used in commercial face recognition systems.
Figure 10: Quantitative comparison of the misdirection success rates of distorted images on three mainstream commercial face recognition systems.
IV-D2 Resisting Face Inpaintings
Another challenge comes from the image inpainting system. A well-trained face inpainting model can recover distorted facial images, rendering defenses ineffective. We evaluate the performance of distorted images against two baseline face inpainting models, namely LBP [54] and GS-SSA [55]. The quantitative results of $L_{2}^{face}$ distance are reported in Table V, and the visualization results are shown in Fig. 11. Although the difference between the repaired distorted image and the forged result is greatly reduced, the proposed method still exhibits optimal defense performance. This is because these face inpainting systems rely heavily on undistorted regions of the image, However, we achieve a greater degree of destruction of the entire image texture due to the introduction of feature loss.
TABLE V: Quantitative results of the image inpainting on distorted images. The best result is marked in bold, while the sub-optimal result is marked with an underline.
  Face inpaintings Methods StarGAN AGGAN FPGAN RelGAN HiSD
LBP [54] Disrupting[10] 0.664 0.114 0.156 0.063 0.053
PG[39] 0.095 0.070 0.092 0.055 0.059
CMUA[12] 0.309 0.089 0.093 0.162 0.072
IAP[4] 0.303 0.094 0.266 0.159 0.080
Ours (S-I) 0.259 0.141 0.335 0.197 0.084
Ours (S-II) 0.387 0.140 0.340 0.214 0.075
  GS-SSA [55] Disrupting[10] 0.772 0.081 0.106 0.028 0.017
PG[39] 0.093 0.033 0.054 0.019 0.024
CMUA[12] 0.437 0.053 0.064 0.148 0.044
IAP[4] 0.318 0.065 0.285 0.134 0.048
Ours (S-I) 0.296 0.108 0.328 0.143 0.059
Ours (S-II) 0.478 0.110 0.374 0.149 0.056
Figure 11: Visual examples of the image inpainting on distorted images. The original and fake images are provided on the right side for reference.
IV-D3 Performance in Gray-box Scenarios
The performance of the proposed method in gray-box scenarios is also evaluated. Gray box scenarios are defined where the model type is known but the internal parameters are not accessible. We train a StarGAN [21] and RelGAN [24] as the target gray-box model respectively. As shown in Table VI, to some extent, ID-guard still maintains the most powerful defense capabilities. Furthermore, it can be found that compared to competing algorithms, ID-guard has the smallest performance degradation. We attribute this to the designed feature loss, which destroys model feature extraction with cross-model consistency [17].
TABLE VI: Quantitative results in gray box scenarios. The best result is marked in bold, while the sub-optimal result is marked with an underline.
  Methods StarGAN RelGAN
[…]
Disrupting[10] 0.187 0.517 0.234 0.201 0.161 0.832
PG[39] 0.094 0.630 0.068 0.044 0.619 0.008
CMUA[12] 0.090 0.798 0.035 0.041 0.719 0.008
IAP[4] 0.106 0.560 0.122 0.114 0.291 0.653
Ours (S-I) 0.129 0.363 0.584 0.107 0.294 0.693
Ours (S-II) 0.109 0.351 0.532 0.124 0.111 0.890
indicates the facial manipulation model in the gray-box setting.
IV-D4 Robustness in Real Social Scenarios
In real scenarios, users often upload perturbed images to social applications to share their lives. However, various lossy operations on the transmission channel can destroy the effectiveness of the perturbation. In this section, we first evaluate the robustness of competing algorithms on Facebook and WeChat. Next, we incorporate the robustness strategy in [18] into our generator training, which can be viewed as a downstream task of this work. Quantitative results are shown in Table VII. Even without robustness training, our method still reports the best robustness compared to the baseline. The reason may be that our adversarial perturbation is not averaged over the image, but is constrained to focus on faces, which makes it harder to neutralize. As shown in Fig. 12, when the robust training strategy is integrated, the anti-lossy operation performance of generated perturbations is greatly improved. This demonstrates the flexibility of the proposed ID-guard framework to effectively integrate with progressive strategies in the community.
TABLE VII: Quantitative results on the robustness of adversarial perturbations in real scenarios. The adversarial images are uploaded to designated social applications and then downloaded. The best result is marked in bold. “w/” and “w/o” indicate whether robust training is used or not, respectively.
[…]
Facebook Disrupting[10] 0.023 0.760 0.022
PG[39] 0.012 0.744 0.009
CMUA[12] 0.003 0.761 0.001
IAP[4] 0.012 0.749 0.016
Ours (S-II) w/o 0.041 0.708 0.112
Ours (S-II) w/ 0.065 0.677 0.306
  WeChat Disrupting[10] 0.010 0.762 0.006
PG[39] 0.004 0.747 0.003
CMUA[12] 0.002 0.741 0.001
IAP[4] 0.004 0.746 0.001
Ours (S-II) w/o 0.017 0.739 0.024
Ours (S-II) w/ 0.052 0.702 0.134
Figure 12: Visual examples of defense effects in real scenarios. The transmission channel is Facebook and the StarGAN is chosen as the target model. “w/” and “w/o” indicate whether robust training is used or not, respectively.
IV-E Further Discussion
IV-E1 How Weights Dynamically Change?
In the proposed ID-guard framework, the dynamic weight strategy is very important, which directly affects the training process and the balance of attack losses for different facial manipulations. Here, to explore its mechanism in depth, we record the dynamic changes of the weight set $\mathcal{S}_{\lambda}=\left\{\lambda_{1},\lambda_{2},\dots,\lambda_{N}\right\}$ of attack losses in an epoch training, as shown in Fig. 13. For strategy I, we use the MGDA algorithm to solve the set of weights automatically in each iteration. It can be found that face manipulation models with strong robustness, such as HiSD [25] and RelGAN [24], tend to be assigned larger weights to strengthen the attack against them. In addition, due to the lack of prior knowledge guidance, the weight change of strategy I is more affected by the current state of the generator, and thus fluctuates more significantly. For strategy II, in the initial stage, there is a large difference between the weights. With the constraints of the prior weights, the allocation of each dynamic weight stabilizes to obtain a balanced performance.
(a) Strategy I
(b) Strategy II
Figure 13: The changing trend of the weight of attack loss with the number of iterations for different facial manipulations when adopting the dynamic weight strategy.
IV-E2 What does the generator learn?
The adversarial perturbations generated by the generator trained under different loss constraints are shown in Fig. 14. The mask loss concentrates rich adversarial information on the facial area of the image, thereby completely distorting the output face. The generator trained using only the identity loss learns to disrupt images at the texture level. Combined with the results in Fig 7, this destruction changes the key feature semantics and visual attributes of the face. Therefore, the standard generator learns to generate adversarial perturbations that cause maximum damage to facial regions and alter the identifiable texture features of the image.
Figure 14: Visual examples of the generated perturbation. Among them, (a) is the original image, (b) is the adversarial image, (c) is the perturbation generated by the generator trained only with mask loss, (d) is the perturbation generated by the generator trained only with identity loss, and (e) is the perturbation generated by the standard generator.
V Conclusion
In this work, we proposed a universal adversarial framework for combating facial manipulation, named ID-guard. To prevent face stigmatization problems caused by unconstrained image distortion, we propose an identity destruction module to eliminate identity semantics. Furthermore, to improve the cross-model performance of generating perturbations, we regard attacking different models as a multi-task learning problem and introduce a dynamic parameter strategy. The proposed method not only effectively resists multiple face manipulations, but also significantly disrupts face identification. In addition, the experiment also demonstrated the possibility of ID-guard in circumventing commercial face recognition systems and image inpaintings. We hope that ID-guard, with its good integration capabilities and application flexibility, can provide the community with an effective solution against facial manipulation.

### v2

@@
 Abstract
-The misuse of deep learning-based facial manipulation poses a potential threat to civil rights.
-To prevent this fraud at its source, proactive defense technology was proposed to disrupt the manipulation process by adding invisible adversarial perturbations into images, making the forged output unconvincing to the observer.
-However, their non-directional disruption of the output may result in the retention of identity information of the person in the image, leading to stigmatization of the individual.
-In this paper, we propose a novel universal framework for combating facial manipulation, called ID-Guard.
-Specifically, this framework requires only a single forward pass of an encoder-decoder network to generate a cross-model universal adversarial perturbation corresponding to a specific facial image.
-To ensure anonymity in manipulated facial images, a novel Identity Destruction Module (IDM) is introduced to destroy the identifiable information in forged faces targetedly.
-Additionally, we optimize the perturbations produced by considering the disruption towards different facial manipulations as a multi-task learning problem and design a dynamic weights strategy to improve cross-model performance.
-The proposed framework reports impressive results in defending against multiple widely used facial manipulations, effectively distorting the identifiable regions in the manipulated facial images.
-In addition, our experiments reveal the ID-Guard’s ability to enable disrupted images to avoid face inpaintings and open-source image recognition systems.
+The misuse of deep learning-based facial manipulation poses a significant threat to civil rights.
+To prevent this fraud at its source, proactive defense has been proposed to disrupt the manipulation process by adding invisible adversarial perturbations into images, making the forged output unconvincing to observers.
+However, the non-specific disruption against the output may lead to the retention of identifiable facial features, potentially resulting in the stigmatization of the individual.
+This paper proposes a universal framework for combating facial manipulation, termed ID-Guard.
+Specifically, this framework operates with a single forward pass of an encoder-decoder network to produce a cross-model transferable adversarial perturbation.
+A novel Identity Destruction Module (IDM) is introduced to degrade identifiable features in forged faces.
+We optimize the perturbation generation by framing the disruption of different facial manipulations as a multi-task learning problem, and a dynamic weight strategy is devised to enhance cross-model performance.
+Experimental results demonstrate that the proposed ID-Guard exhibits strong efficacy in defending against various facial manipulation models, effectively degrading identifiable regions in manipulated images.
+It also enables disrupted images to evade facial inpainting and image recognition systems.
+Additionally, ID-Guard can seamlessly function as a plug-and-play component, integrating with other tasks such as adversarial training.
 Index Terms:
@@
 I Introduction
-The spread of false information in communities has been a longstanding concern, presenting a potential threat to civil rights and social security.
-In recent years, the advancement and deployment of generative deep neural networks (DNNs) have exacerbated this issue, with facial manipulation serving as a notable example.
-This technology enables end-to-end manipulation of facial attributes or identity of images/videos.
-Malicious actors, for instance, exploit forged images to generate and circulate misleading news [1, 2] or perpetrate online fraud [3].
-Although the re-training of these methods is proved challenging due to high computing power requirements and technical thresholds, users can easily download accessible pre-trained models from open-source platforms such as Github  [^1] https://github.com, Hugging Face  [^2] https://huggingface.com, and TensorFlow Hub  [^3] https://www.tensorflow.org to implement the forgery [4].
-This greatly reduces the cost of creating fake examples, thereby expediting the proliferation of false information on social media.
-Therefore, an urgent need exists to develop proactive and efficient defense mechanisms.
+The spread of false information in communities has long been a major concern, posing a potential threat to civil rights and social security.
+The rapid advancement and widespread deployment of generative deep neural networks (DNNs) have further intensified this issue, with facial manipulation emerging as a prominent example.
+This technology enables end-to-end manipulation of facial attributes or identities of images and videos.
+Malicious actors, for instance, leverage forged images to fabricate and disseminate misleading news [7, 8] or engage in online fraud [6].
+Although re-training these models remains challenging due to substantial computational demands and technical barriers, pre-trained models are readily available on open-source platforms such as GitHub  [^1] https://github.com, Hugging Face  [^2] https://huggingface.com, and TensorFlow Hub  [^3] https://www.tensorflow.org, enabling users to easily execute forgeries [9].
+This significantly lowers the barrier to generating fake content, thereby accelerating the spread of misinformation on social media.
+Consequently, there is an urgent need to develop effective and proactive defense mechanisms.
 Figure 1: Illustration of the impact of malicious propagation of facial manipulation samples.
-Fakes will lead to rumor spreading and insufficient distortion of faces by traditional defense methods will cause face stigmatization.
+Fakes will lead to rumors spreading, and insufficient distortion of faces by traditional defense methods will cause face stigmatization.
 Our method disrupts the observer’s identification of the identity in the sample and thus adequately protects the individual’s rights.
-To mitigate the aforementioned threats, substantial research efforts have been directed toward proactive defense mechanisms against facial manipulation in recent times.
-Unlike passive detection methods [5, 6, 7, 8, 9], proactive defense algorithms [10, 11, 12, 4, 13, 14, 15, 16, 17, 18, 19] are crafted to thwart fraudulent activities at their source.
-These algorithms induce visual distortions in facial manipulation outputs by introducing imperceptible adversarial perturbations [20] into face images.
-However, the disruption caused by existing algorithms to manipulated images is often non-directional, leading to the preservation of individuals’ identifiable information in the disrupted images.
-When stubborn malicious users insist on uploading these incompletely disrupted forged images to social media, concerns such as face stigmatization may arise [16].
-As illustrated in Fig.
-1, applying the traditional proactive defense algorithm mentioned above to safeguard photographs of the renowned movie star Jackie Chan still allows ordinary Internet users to discern his identity readily.
-Insufficient distortion of facial images makes the disrupted output a “spoof”, leaving the victim’s reputation and rights inadequately protected.
-Notably, as shown in Fig.
-2, above insufficiently disrupted facial examples also face two potential threats: 1) The residual identifiable information in the samples makes them still likely to be recognized by commercial facial recognition systems.
-The stigmatization problem is aggravated by the fact that some entertainment applications automatically identify and push images of celebrities; 2) Some malicious forgers with rich technical capabilities may restore the fake images that are not seriously distorted to continue to commit fraud.
-Figure 2: Illustration of potential threats to the insufficiently disrupted facial example. Challenges come primarily from commercial face recognition systems and face inpainting algorithms.
-To address the above concerns, in this paper, we propose a proactive defense framework called ID-guard.
-This framework requires only a single forward pass of an image reconstruction network to produce the universal adversarial perturbation to combat multiple open-source facial manipulation algorithms, as shown in Fig.
+In response to these threats, significant research efforts have recently focused on developing proactive defense mechanisms against facial manipulation.
+Unlike passive detection methods [18, 19, 20, 56, 57], proactive defense algorithms [12, 23, 11, 9, 24, 13, 14, 15, 25, 27, 16] are designed to counteract fraudulent activities at their origin.
+However, these distortions suffer from several critical limitations: 1) they fail to completely obscure personally identifiable features, allowing identity-relevant information to persist, as the distortions produced by unconstrained adversarial perturbations appear in indeterminate locations rather than effectively covering the entire face; 2) they introduce random and unstructured artifacts, such as shadows, background distortions, and deformed facial regions, resulting in unnatural and visually unacceptable appearances.
+Consequently, if persistent malicious users insist on uploading these non-specifically disrupted images to social media, and these images are maliciously disseminated, concerns such as facial stigmatization may arise, posing ethical and reputational risks for the individuals depicted [15].
+Especially for public figures, their identities remain discernible even after defense mechanisms are applied.
+For example, as illustrated in Fig.
+1, when the above-mentioned traditional proactive defense is applied to protect a photo of the famous movie star Jackie Chan, his identity remains recognizable from the stigmatized forged output.
+Furthermore, as illustrated in Fig.
+2, inadequately disrupted facial images are also exposed to two potential threats: 1) The remaining identifiable information in these images increases the likelihood of their recognition by commercial facial recognition systems.
+This exacerbates the issue of stigmatization, as certain entertainment applications automatically detect and promote images of celebrities; 2) Technically adept malicious forgers may restore forged images that have not been significantly distorted by facial inpainting, enabling continued fraudulent activities.
+Figure 2: Illustration of potential threats to the insufficiently disrupted facial example. Challenges come primarily from commercial face recognition systems and facial inpainting algorithms.
+Figure 3: The publicly available pre-trained models can be easily downloaded from open-source platforms to implement forgeries.
+For a given image, the proposed ID-Guard can generate transferable perturbations for defense against multiple open-source facial manipulations through a single forward propagation of an image reconstruction network.
+To address the above concerns, in this paper, we propose a proactive defense framework, ID-Guard.
+The framework generates transferable adversarial perturbations through a single forward pass of an image reconstruction network, effectively countering multiple open-source facial manipulation algorithms, as shown in Fig.
 3.
-To destroy the identifiable semantic information in forged images and prevent spoofing of individuals, a novel Identity Destruction Module (IDM) is introduced.
-IDM enables the generated protective perturbations to focus on disrupting key regions of the individual’s face in an image to prevent an observer from recognizing his/her identity.
-In addition, the universality of the adversarial perturbations is crucial in practical application scenarios [12] due to the facial manipulation methods used by the forger are unknown and uncontrollable.
-To improve the cross-model universality of perturbations, a dynamic weighting strategy is proposed during the training of the perturbation generator.
-Specifically, the robustness of different facial manipulations against attacks varies due to differences in model structure and complexity.
-Simply treating the adversarial loss of attacking different models as equally weighted when training the perturbation generator will result in the produced perturbations being biased towards facial manipulations that are easy to attack.
-Therefore, we consider the attacks against different models as a multi-task learning problem, which enables the loss weights to be dynamically adjusted during the training process to achieve well-balanced cross-model performance.
-A gradient prior perturbation strategy is also introduced to guide the perturbation generator to improve the stability of training and accelerate convergence.
-As expected and will be verified by experiments, the proposed ID-guard can effectively distort recognizable regions in facial images forged by various open-source manipulation algorithms, thereby preventing observers and face recognition systems from recognizing their identities and circumventing face inpaintings.
-In summary, the contributions of our work are summarized as follows:
+To eliminate identifiable semantic information in forged images and prevent identity spoofing, a novel Identity Destruction Module (IDM) is incorporated.
+The IDM disrupts identity-related features in a structured manner rather than introducing random or uncontrolled distortions, thereby ensuring that the individual’s identity remains unrecognizable.
+The transferability of adversarial perturbations is crucial in practical applications, as the facial manipulation methods employed by forgers are often unknown and uncontrollable.
+To address this, a dynamic weight strategy is introduced during the training of the perturbation generator.
+Specifically, the robustness of different facial manipulation models varies due to differences in model architecture and method design.
+Assigning equal weights to the adversarial loss of different models during perturbation generator training may cause the produced perturbations to be biased toward easily attacked facial manipulations, thereby degrading overall performance.
+Thus, adversarial attacks on different models are formulated as a multi-task learning problem, allowing the loss weights to be dynamically adjusted during training to achieve well-balanced cross-model performance.
+Additionally, a gradient prior perturbation strategy is introduced to enhance training stability and accelerate convergence.
+As anticipated and confirmed through experiments, the proposed ID-Guard effectively distorts identifiable regions in facial images manipulated by various open-source models, thereby preventing observers and face recognition systems from identifying individuals and bypassing facial inpainting techniques.
+Additionally, we demonstrate that ID-Guard can function as a plug-and-play module in adversarial training for facial manipulation models, thereby enhancing their adversarial robustness.
+In summary, the contributions of this work are summarized as follows:
   1. 1.
-    We propose a novel general adversarial perturbation generation framework to prevent facial manipulation from stigmatizing individuals, called ID-guard.
-Only a single forward pass of the generator is required to quickly produce perturbations that can disrupt various facial manipulations.
+    We propose a novel general adversarial perturbation generation framework, termed ID-Guard, to prevent facial manipulation from stigmatizing individuals.
+A single forward pass of the generator suffices to produce perturbations capable of disrupting various facial manipulations.
+Moreover, this framework can be leveraged for seamless integration with other tasks.
   2. 2.
-    To fully disrupt the manipulation results so that the identities of individuals in the images cannot be identified, an Identity Destruction Module is proposed to guide the generated perturbations to focus on distorting identity-related semantic information.
-This module can further alleviate concerns about commercial face recognition systems and image inpainting algorithms.
+    To ensure complete disruption of manipulated images, preventing the identification of individuals, an Identity Destruction Module (IDM) is introduced.
+The IDM guides the generated perturbations to target identity-related semantic features, thereby mitigating concerns regarding commercial face recognition systems and image inpainting algorithms.
   3. 3.
-    To improve the universality of the generated adversarial perturbations, we implement an attack on multiple facial manipulations by solving a multi-task learning problem and designing a dynamic weighting strategy.
-To improve the stability of the generator, we introduce a gradient prior perturbation strategy.
-The remainder of this paper is organized as follows: Related works on Deepfakes and proactive defense are reviewed in Section II.
+    To improve the transferability of the generated adversarial perturbations, we implement attacks against multiple facial manipulations by solving a multi-task learning problem and designing a dynamic weight strategy.
+To improve the stability of the generator, a gradient prior perturbation strategy is introduced to improve the generator’s stability.
+The remainder of this paper is organized as follows: Related works on facial manipulation and proactive defense are reviewed in Section II.
 Section III presents the details of our method.
 The experimental results and analysis are provided in Section IV, followed by the conclusion in Section V.
-Figure 3: The accessible pre-trained models can be easily downloaded from open-source platforms to implement forgeries.
-For a given image, the proposed ID-guard can generate universal perturbations for defense against multiple open-source facial manipulations through a single forward propagation of an image reconstruction network.
 II Related Works
 II-A Facial Manipulation
-Facial manipulation aims at constrained modification of the face of a given image/video to present the desired visual content, such as identity, expression, age, hair color, etc.
-of the character.
-In recent years, thanks to the great success of Generative Adversarial Networks (GANs) in image synthesis, many GAN-based algorithms [21, 22, 23, 24, 25] with different designs and constraints have been proposed to solve facial manipulation [26].
-Some contributors choose to open source their work on the public platform, including pre-training models and running scripts, which greatly lowers the threshold for users to produce high-quality and high-fidelity fake images/videos.
-II-B Facial Manipulation Disruption
-Some recent works have achieved proactive defense against facial manipulation by injecting adversarial perturbations into the image.
-Ruiz et al.[10] and Yeh et al.[11] achieved disruption of the facial manipulation by deriving gradient-based adversarial perturbations on the target model.
-Works including [18] and [19] have significantly improved the robustness of adversarial perturbations for protecting personal images.
-However, the common drawbacks of these methods are higher computational overhead and lower work efficiency, due to they involve solving the model-specific perturbation for each image.
-Methods represented by [12, 4, 17] and [15] have studied cross-model universal perturbations, which to some extent alleviate overhead issues and practicality limitations.
-However, they ignore the fact that different facial manipulation models have differences in robustness and gradient direction, and thus the performance of the produced generic perturbations is uneven across models.
-Furthermore, as discussed above, these algorithms do not consider the problem of face stigmatization due to unconstrained disruptions.
+Facial manipulation refers to the controlled modification of facial attributes in a given image or video to generate the desired visual content, including identity, expression, age, and hair color.
+In recent years, leveraging the remarkable success of Generative Adversarial Networks (GANs) in image synthesis, numerous GAN-based algorithms [1, 2, 3, 4, 5] featuring diverse architectures and constraints have been developed to facilitate facial manipulation [22].
+Some researchers have opted to release their work as open-source on public platforms, providing pre-trained models and executable scripts, thereby significantly reducing the technical barriers for users to generate high-quality, high-fidelity fake images and videos.
+II-B Proactive Defense against Facial Manipulation
+From a defense objective standpoint, generalized proactive defense methods can be divided into two categories: proactive forensics and proactive disruption.
+Proactive forensics involves embedding imperceptible watermarks or traceable markers into multimedia content to facilitate the identification of manipulated samples.
+Using these embedded elements, defenders can verify content authenticity and trace the origins of facial manipulations.
+In contrast, proactive disruption seeks to degrade the quality of facial manipulation outputs through the injection of adversarial perturbations, thereby misleading the generative model.
+By distorting the generated results, proactive disruption effectively diminishes the realism and credibility of forged images.
+II-B1 Proactive Forensics
+Proactive forensic techniques aim to embed identifiable patterns into images to facilitate fake detection and manipulation provenance tracking.
+Early approaches, such as FaceGuard [58] and Faketagger[59], detected forged examples by embedding watermarks into real images and verifying their integrity upon retrieval, but lacked structured tracking mechanisms.
+To provide identity source tracking, Zhao et al.
+[63] embed watermarks as anti-Deepfake labels into facial identity features, enabling fake detection by verifying the presence of the label.
+These works mainly focus on authenticity detection, but cannot pinpoint the tampered region.
+To improve detection and localization, Asnani et al proposed a proactive embedding framework [60], later refining it into MALP [61], which integrates attention mechanisms to achieve fine-grained manipulation localization.
+PADL [62] further enhanced robustness by combining perturbation-based defenses with detection and localization strategies.
+Zhao et al.
+[64] embedded a semi-fragile watermark in the original image.
+Once counterfeited, the tampered regions can be located by comparing retrieved and original watermarks.
+The limitation of proactive forensics is that it preserves the integrity of forged samples, whereas proactive disruption directly degrades or nullifies the forgery.
+Therefore, in this paper, we focus on proactive disruption methods.
+II-B2 Proactive Disruption
+Recent studies have explored proactive disruption against facial manipulation by injecting adversarial perturbations into images.
+Ruiz et al.[12] and Yeh et al.[23] disrupted facial manipulation by deriving gradient-based adversarial perturbations on target models.
+Works including [27] and [16] have significantly improved the robustness of adversarial perturbations in protecting personal images.
+However, due to structural and design differences across various facial manipulation models, adversarial perturbations crafted for a specific model often exhibit poor transferability to other models, limiting their defense applicability.
+To address this, works such as [13] and [10] generated adversarial perturbations by attacking a surrogate model and transferred them to an inaccessible model.
+However, the significant structural differences between face manipulation models limit their effectiveness.
+A more widely adopted approach is model ensembling.
+Representative methods such as [11, 9, 25] and [14] have explored cross-model transferable adversarial perturbations based on this paradigm, which enhance defense effectiveness against various facial manipulation models to a certain extent.
+However, these approaches overlook the fact that different facial manipulation models exhibit variations in adversarial robustness and gradient optimization.
+As a result, the effectiveness of the generated transferable perturbations is inconsistent across models, leading to an overall decline in defense performance.
+This is one of the issues that this paper focuses on.
+Additionally, as noted earlier, these methods do not account for the issue of facial stigmatization caused by unconstrained perturbations.
 Zhai et al.
-[16] proposed a method to embed specific warning patterns on the generated fake images to solve this problem.
+[15] addressed this problem by embedding specific warning patterns into generated fake images.
 Unlike them, the proposed ID-Guard directly distorts the facial recognition area of fake images.
@@
 Some heuristic works [31, 32, 34, 33, 35] measured the difficulty of a task based on the order of magnitude or change rate of the loss value, and then dynamically adjusted the weights of different tasks to obtain balanced performance.
-In this work, we further explore the potential of integrating multi-task learning strategies into across-model universal perturbation generation.
+In this work, we further explore the potential of integrating multi-task learning strategies into across-model transferable perturbation generation.
 III Methodology
 In this section, the specific design and implementation details of the proposed ID-Guard framework are elaborated.
-For clarity, we first introduce the overview of the framework and the definition of notations.
-III-A Preliminaries
-III-A1 Adversarial Attacks against Facial Manipulations
-Given a natural image $x\in\mathbb{R}^{3\times H\times W}$, the pre-trained facial manipulation model $\mathcal{M}$ translates it into a forgery $y=\mathcal{M}(x)$.
-The goal of the defender is to find a small adversarial perturbation $\delta$ that makes the facial manipulation model fail to manipulate the perturbed adversarial image $x_{adv}=x+\delta$.
-The solution to the $\delta$ can be formulated as a maximization optimization problem as follows
-$\begin{split}&\max_{\eta}\mathcal{D}(\mathcal{M}(x),\mathcal{M}(x+\delta)),\\ &s.t.\quad\|\delta\|_{\infty}\leq\epsilon,\end{split}$ (1)
-where $\mathcal{D}$ is the metric of the distance between images and its design is our focus and will be introduced in subsequent sections.
-$\epsilon$ is the infinite norm bound used to restrict the perturbation.
-The existing mainstream adversarial perturbation derivation patterns can be roughly divided into two categories: gradient-based methods [34, 35] and generation-based methods[36, 37].
-Compared to the former, the latter is more flexible and faster, so we choose the generation-based approach to drive the perturbation derivation of the proposed ID-guard.
-Specifically, we train a perturbation generator $\mathcal{G}$ with the following optimization objective to produce the desired adversarial perturbation:
-$\begin{split}&\max_{{\theta}_{\mathcal{G}}}\mathbb{E}(\mathcal{D}(\mathcal{M}(x),\mathcal{M}(x+\mathcal{G}(x)))),\\ &s.t.\quad\|\mathcal{G}(x)\|_{\infty}\leq\epsilon\end{split}$ (2)
-$\mathcal{G}$ is designed as a Resnet architecture [38], and the advantages of this design will be discussed in subsequent experiments.
-III-A2 Cross-Model Universal Perturbations
-We hope that the adversarial perturbation generated for a certain image can be effective for a set of open-source pre-trained facial manipulation models $\mathcal{S}_{\mathcal{M}}=\left\{\mathcal{M}_{1},\mathcal{M}_{2},\dots,\mathcal{M}_{N}\right\}$.
-Hence, the optimization objective in Eq.
-(2) can be rewritten as:
+For clarity, we first introduce an overview of the framework and a definition of notation.
+III-A Overview
+III-A1 Facial Manipulation
+Facial manipulation can be regarded as an image translation task, aiming to transform a given original example into a target manipulated example.
+Specifically, given an original face image $x\in\mathbb{R}^{3\times H\times W}$, the facial manipulation model $\mathcal{M}$ leverages the specified target attribute or identity $a$ to map it to a forged image $y$.
+The process is formulated as:
+$\begin{split}y=\mathcal{M}(x,a).\end{split}$ (1)
+Since end-to-end facial manipulation models employ different methods for embedding attributes or identity information, we simplify the manipulation process as $y=\mathcal{M}(x)$.
+where the face in the manipulated image $y$ retains the identity of the person in the original image $x$ while exhibiting the specified features of attribute $a$.
+III-A2 Adversarial Perturbation against Manipulation
+To proactively disrupt facial manipulation, the defender’s objective is to generate an imperceptible adversarial perturbation $\delta$ such that the target facial manipulation model fails to map the adversarial image $x_{adv}=x+\delta$ to an acceptable manipulated output.
+The optimization of $\delta$ can be formulated as a maximization problem:
+$\begin{split}&\max_{\delta}\mathcal{D}(\mathcal{M}(x),\mathcal{M}(x+\delta)),\\ &s.t.\quad\|\delta\|_{\infty}\leq\epsilon,\end{split}$ (2)
+where $\epsilon$ is the infinite norm bound used to restrict the perturbation, and $\mathcal{D}$ is the distance metric between the original forged image $\mathcal{M}(x)$ and disrupted forged image $\mathcal{M}(x+\delta)$.
+Some existing works [12, 23] employ gradient-based adversarial attack algorithms[36, 37] to generate perturbations.
+However, these methods require multiple iterations for each perturbation generation, leading to high computational overhead.
+A more efficient approach is to train a perturbation generator $\mathcal{G}$, which can produce adversarial perturbations with a single forward pass during inference, i.e., $\delta=\mathcal{G}(x)$.
+The optimization problem for training $\mathcal{G}$ is defined as follows:
+$\begin{split}&\max_{{\theta}_{\mathcal{G}}}\mathbb{E}(\mathcal{D}(\mathcal{M}(x),\mathcal{M}(x+\mathcal{G}(x)))),\\ &s.t.\quad\|\mathcal{G}(x)\|_{\infty}\leq\epsilon,\end{split}$ (3)
+where ${\theta}_{\mathcal{G}}$ is the parameter of the perturbation generator.
+On one hand, previous state-of-the-art methods [12, 23, 10, 11] typically apply the Mean Squared Error (MSE) loss as a proxy for the distance metric $\mathcal{D}$.
+However, this approach often results in unstructured distortions in the manipulated image while preserving identifiable facial features, leading to facial stigmatization.
+On the other hand, the perturbations generated by these methods generally lack transferability and can only disrupt a single target model $\mathcal{M}$.
+III-A3 ID-Guard Framework
+To address the aforementioned issues, in this paper, we propose the ID-Guard framework, which aims to train a generator $\mathcal{G}$ with Resnet[51] architecture capable of producing the cross-model transferable adversarial perturbation, as shown in Fig.
+4.
+This generator produces customized perturbations for given images, effectively defending against a set of pre-trained facial manipulation models $\mathcal{S}_{\mathcal{M}}=\left\{\mathcal{M}_{1},\mathcal{M}_{2},\dots,\mathcal{M}_{N}\right\}$.
+Following the model ensemble paradigm, for all accessible models $\mathcal{M}_{k}$, Eq.(3) in our proposed framework can be rewritten as:
 $\begin{split}&\max_{{\theta}_{\mathcal{G}}}\sum_{k=1}^{N}\lambda_{k}\mathbb{E}(\mathcal{D}(\mathcal{M}_{k}(x),\mathcal{M}_{k}(x+\mathcal{G}(x)))),\\ &s.t.\quad\|\mathcal{G}(x)\|_{\infty}\leq\epsilon\end{split}$ (4)
-where $\mathcal{S}_{\lambda}=\left\{\lambda_{1},\lambda_{2},\dots,\lambda_{N}\right\}\in\mathbb{R}$ is a set of weights got based on the robustness of each target model.
-Solving the effective $\mathcal{S}_{\lambda}$ is one of the key tasks of this work so that the generated adversarial perturbations have a balanced defense performance against different facial manipulations.
-This part will be introduced in detail in Section III-C.
+where $\mathcal{S}_{\lambda}=\left\{\lambda_{1},\lambda_{2},\dots,\lambda_{N}\right\}$ represents a set of weights used to balance the adversarial loss for different models during training.
+Determining these weights is one of our key tasks.
+To overcome the adversarial gradient discrepancies caused by significant structural differences among facial manipulation models, we propose a dynamic weight strategy to achieve balanced cross-model defense performance, which will be introduced in Section III-C.
+Additionally, to address the challenge of facial stigmatization caused by non-specific constraint distortions, we propose an Identity Destruction Module (IDM) to compute the distance metric $\mathcal{D}$, which will be detailed in Section III-B.
+Moreover, to incorporate structural adversarial gradient priors during the training of $\mathcal{G}$, we introduce a Gradient Prior Strategy, which is discussed in Section III-D.
 Figure 4: Illustration of the proposed ID-Guard framework.
 The perturbation generator takes a natural image $x$ as input and requires only one forward propagation to generate a cross-model adversarial perturbation dedicated to the input face that can be used to defend against multiple facial manipulations.
-In the training phase, ID-guard consists of three modules, including the identity destruction module, the dynamic weighting strategy, and the gradient prior perturbation strategy.
+In the training phase, ID-Guard consists of three modules, including the Identity Destruction Module, the dynamic weight strategy, and the gradient prior perturbation strategy.
 The notation descriptions are shown in the upper right corner for reference.
 III-B Identity Destruction Module
-As mentioned above, the design of the distance metric $\mathcal{D}$ is very critical.
-Traditional methods [10, 11, 39, 12] generally choose Mean Squared Error (MSE) as a proxy for distance measurement.
-However, this approach causes non-directional distortion of forged images and fails to ensure that the identity of the distorted face cannot be recognized.
-As an improvement, we introduce the Identity Destruction Module (IDM).
+To clarify the algorithmic design, we first introduce the Identity Destruction Module (IDM).
+Based on this module, we then provide a complete definition of the training loss function for the perturbation generator $\mathcal{G}$.
 As shown in Fig.
 4, the IDM consists of three sub-modules, which will be introduced separately next.
-III-B1 Mask Loss
+III-B1 Mask Constrained Loss
 First, we consider using face masks to limit the regions of image distortion by adversarial perturbations.
-The designed mask is two-fold: 1) The binary mask is used to restrict image distortion to areas of facial components including the eyes, nose, mouth, and eyebrows, which are proven to play an important role in identity recognition by the human eye [40, 41, 42]; 2) The heatmap mask weights the face distortion loss at the pixel level, making the perturbation pay more attention to the important feature areas of the face.
-In this work, the heatmap of each image is obtained by solving Grad-SAM [43] on VGGFace [44].
+The designed mask is two-fold: 1) The binary mask is used to restrict image distortion to areas of facial components including the eyes, nose, mouth, and eyebrows, which are proven to play an important role in identity recognition by human eyes [42, 40, 41]; 2) The heatmap mask weights the face distortion loss at the pixel level, making the perturbation pay more attention to the important feature areas of the face.
+In this work, the heatmap of each image is obtained by solving Grad-cam [45] on VGGFace [44].
 This design will also facilitate distorted images against commercial facial recognition systems.
@@
 $\mathcal{L}_{k}^{mask\_hm}=\|\mathcal{M}_{k}(x)\odot m^{hm}-\mathcal{M}_{k}(x+\mathcal{G}(x))\odot m^{hm}\|_{2}$ (6)
-where $m^{bin}$ and $m^{hm}$ denote the binary mask and heatmap mask of the natural image $x$, respectively.
+where $m^{bin}$ and $m^{hm}$ denote the binary mask and heatmap mask of the original image $x$, respectively.
 Note that these masks are only computed during the perturbation generator training stage to constrain the distortion region and are not used in the inference process.
 $\odot$ indicates the element-wise multiplication.
-III-B2 Identity Loss
-In addition to pixel-level constraints, further consideration is given to maximizing the identity similarity of the manipulation results on the original and adversarial images.
-We achieve this by maximizing the cosine distance between their identity embeddings:
+III-B2 Identity Consistency Loss
+In addition to pixel-level constraints, we also consider maximizing the identity discrepancy between the forgery outputs of the facial manipulation model $\mathcal{M}_{k}$ for the original image and the adversarial image, i.e., maximizing the discrepancy between the original forged image $\mathcal{M}_{k}(x)$ and the disrupted forged image $\mathcal{M}_{k}(x_{adv})$.
+This is achieved by maximizing the cosine distance between their identity embeddings:
 $\mathcal{L}_{k}^{ID}=\cos(ID(\mathcal{M}_{k}(x)),ID(\mathcal{M}_{k}(x+\mathcal{G}(x))))$ (7)
 where $ID(\cdot)$ denotes the Arcface net [46] designed to extract high-quality features from facial images and embed them into a low-dimensional space where the distance between different embeddings corresponds to the similarity between faces.
-III-B3 Feature Loss
-Learning from [17] that the process of attribute or identity embedding for end-to-end facial manipulation varies, but the feature extraction process is similar.
+III-B3 Feature Confusion Loss
+It is crucial to exploit the commonalities between different facial manipulation models during perturbation generation.
+Learning from [25] that while the attribute embedding processes in end-to-end facial manipulation models vary, their feature extraction processes share similarities.
 In addition, feature-level perturbations can retain their effective components in network transmission to a greater extent [15].
-Therefore, a feature loss that enables the generator to focus on destroying feature-level faces is incorporated into the IDM to improve the effectiveness and transferability of produced perturbations:
+Therefore, a feature confusion loss that enables the generator to focus on destroying feature-level faces is incorporated into the IDM to improve the effectiveness and transferability of produced perturbations:
 $\mathcal{L}_{k}^{feat}=\|E_{k}(x)\odot LI(m^{bin})-E_{k}(x+\mathcal{G}(x))\odot LI(m^{bin})\|_{2}$ (8)
@@
 $LI(\cdot)$ represents the linear interpolation operation to make the binary mask and the extracted feature map consistent in image size.
-In summary, given a pre-trained facial manipulation model $\mathcal{M}_{k}\in\mathcal{S}_{\mathcal{M}}$, the attacking loss against it can be formulated as:
-$\mathcal{L}_{k}=\mathcal{L}_{k}^{mask\_bin}+\mathcal{L}_{k}^{mask\_hm}+\mathcal{L}_{k}^{ID}+\mathcal{L}_{k}^{feat}$ (8)
-The loss function for training the perturbation generator $\mathcal{G}$ is a linear combination of attacking losses of facial manipulation models:
-$\mathcal{L}=\lambda_{1}\cdot\mathcal{L}_{1}+\lambda_{2}\cdot\mathcal{L}_{2}+\dots+\lambda_{N}\cdot\mathcal{L}_{N}$ (9)
+In summary, given a pre-trained facial manipulation model $\mathcal{M}_{k}\in\mathcal{S}_{\mathcal{M}}$, the adversarial loss against it can be formulated as:
+$\mathcal{L}_{k}^{adv}=\mathcal{L}_{k}^{mask\_bin}+\mathcal{L}_{k}^{mask\_hm}+\mathcal{L}_{k}^{ID}+\mathcal{L}_{k}^{feat}$ (9)
+The loss function for training the perturbation generator $\mathcal{G}$ is a linear combination of adversarial losses of facial manipulation models:
+$\mathcal{L}^{adv}=\lambda_{1}\cdot\mathcal{L}_{1}^{adv}+\lambda_{2}\cdot\mathcal{L}_{2}^{adv}+\dots+\lambda_{N}\cdot\mathcal{L}_{N}^{adv}$ (10)
 The definition of $\lambda_{k}$ is shown in Eq. (4), and the use of the proposed dynamic weight strategy to solve it will be introduced next.
 III-C Dynamic Weight Strategy
-We regard the attack against different facial manipulation models in Eq.
-(9) as a multi-task learning problem to solve the weight set $\mathcal{S}_{\lambda}$.
-As shown in Fig 4, the proposed dynamic weighting strategy is two-fold.
-First, a classical but effective multi-task learning strategy, i.e., MGDA [28], is integrated into the proposed ID-guard framework in order to dynamically adjust the weights to optimize the gradients of multiple tasks during the generator training phase.
-Specifically, we follow [27]  [^4] https://github.com/isl-org/MultiObjectiveOptimization to solve this multi-objective optimization problem and use it as a baseline dynamic weighting strategy (noted as S-I).
-On the other hand, we introduce a strategy that combines prior weights and dynamic weights (noted as S-II). For the $\lambda_{k}$ of the attack loss in Eq. (9), refine it to
+We regard the attacks against different facial manipulation models in Eq.
+(10) as a multi-task learning problem to optimize the weight set $\mathcal{S}_{\lambda}$, ensuring that the generated perturbations achieve more balanced cross-model adversarial effectiveness.
+As shown in Fig.
+4, the proposed dynamic weight strategy is two-fold.
+III-C1 MGDA-based
+First, we integrate the Multiple Gradient Descent Algorithm (MGDA) [30] into the proposed ID-Guard.
+The goal of MGDA is to find a set of weights $\lambda_{k}$ that balances the trade-offs between tasks while respecting the constraints of Pareto optimality.
+It can be formulated as the following quadratic programming problem:
+$\begin{split}&\min_{\lambda\geq 0}\quad\left\|\sum_{k=1}^{N}\lambda_{k}\mathcal{L}_{k}^{adv}\right\|^{2},\\ &s.t.\quad\sum_{k=1}^{N}\lambda_{i}=1\end{split}$ (11)
+In this paper, we employ the Frank-Wolfe method [29] to solve the optimization problem.
+The core idea is to combine gradients from multiple tasks into a suitable descent direction that minimizes the losses across all tasks.
+Specifically, we follow the implementation outlined in [29]  [^4] https://github.com/isl-org/MultiObjectiveOptimization.
+This dynamic weight strategy, based on MGDA, is used as a baseline, referred to as Strategy I (S-I).
+III-C2 KPI-based
+Additionally, a dynamic weight strategy based on Key Performance Indicators (KPI) is introduced, denoted as Strategy II (S-II).
+Each dynamic weight $\lambda_{k}$ is refined into a prior weight and a learnable weight, as follows:
 $\lambda_{k}^{t}=\alpha_{k}\times\beta_{k}^{t}$ (12)
-where $t$ indexes the iteration steps, $\alpha_{k}=10^{n}\,(n\in\mathbb{Z})$ is a prior order of magnitude weight for the attack loss, which is determined based on the adversarial robustness of the corresponding facial manipulation model.
-Here, a heuristic is used to quantify the adversarial robustness: on a mini-batch of images, the equivalent setting PGD [35] is used to derive a gradient-based adversarial attack against facial manipulation models, and the $L_{2}$ distance of the attack is calculated.
-$\beta_{k}^{t}$ is dynamically adjusted during the training stage.
-Inspired by [31], we would like to make the harder-to-learn task, i.e., the harder-to-attack facial manipulation model, have higher weights during training.
-Hence, $\beta_{k}^{t}$ is solved as:
-$\beta_{k}^{t}=-(1-\mathcal{K}_{k}^{t})\log\mathcal{K}_{k}^{t}$ (11)
-where $\mathcal{K}_{k}$ represents the KPI for $\mathcal{M}_{k}$, which is an indicator of attack performance.
-KPI is inversely proportional to the training difficulty of the task, that is, the higher the KPI, the easier the task is to learn.
-Here we choose $L_{2}$ distance as the proxy of KPI.
-Intuitively, tasks with high KPIs are easier to learn and therefore become less weighted; conversely, tasks that are difficult to learn become more weighted.
-This combined design improves the stability of dynamic weights to produce adversarial perturbations with more balanced performance.
+where $t$ indexes the iteration steps, $\alpha_{k}=10^{n}\,(n\in\mathbb{Z})$ is the prior weight representing the magnitude of the attack loss, which reflects the adversarial robustness of different forged models, and $\beta_{k}^{t}$ is the learnable weight updated adaptively based on defense performance during each iteration.
+To set the prior weight $\alpha_{k}$ reasonably, a small-sample heuristic method is used to quantify the prior adversarial robustness of different facial manipulation models.
+Specifically, a small batch of face images is randomly selected, and slight attacks are applied to each model using PGD [37] with the same settings.
+The magnitude of the $L_{2}$ distance is used to determine the value of the prior weight.
+This approach provides a reasonable initialization of weights in the early stages of training, balancing the contribution of adversarial losses across different models.
+For the learnable weight $\beta_{k}^{t}$, inspired by the study [34], we assign higher loss weights to more difficult-to-learn tasks, i.e., facial manipulation models that are harder to disrupt during training.
+The $\beta_{k}^{t}$ is computed as follows:
+$\begin{split}&\beta_{k}^{t+1}=-(1-\mathcal{K}_{k}^{t})\log\mathcal{K}_{k}^{t},\\ &\beta_{k}^{0}=1\end{split}$ (13)
+where $\mathcal{K}_{k}^{t}$ is the KPI representing the attack difficulty of model $\mathcal{M}_{k}$ at iteration $t$.
+A higher KPI indicates that the model is easier to attack, so a smaller weight $\beta_{k}^{t}$ is assigned, while a lower KPI $\mathcal{K}_{k}$ results in a larger weight to enhance the adversarial impact of the generated perturbation against the model $\mathcal{M}_{k}$.
+Thus, the KPI value is inversely proportional to the adversarial loss weight.
+A negative correlation function of $\mathcal{K}_{k}^{t}$ is used to calculate the $\beta_{k}^{t}$.
+For convenience, we use the proposed $\mathcal{L}_{k}^{mask\_hm}$ to compute the distance between the real and perturbed fake results as a proxy for KPI, with its range truncated to $(0,1]$ to ensure the monotonicity of the function.
+Unlike the MGDA-based dynamic weight strategy, which requires additional computation in each iteration, the KPI-based strategy directly uses the loss function value calculated in the previous iteration as the KPI to obtain the learnable weight for this iteration.
+This combination design is more efficient and enhances the stability of the dynamic weight.
 III-D Gradient Prior Perturbation
@@
 where $\mathrm{cat}(\cdot)$ denotes channel-wise concatenation, i.e., $\mathrm{cat}(x,\delta_{p})\in\mathbb{R}^{6\times H\times W}$.
-The intuition behind this design is that the gradient prior perturbation can provide the generator with rich prior gradient and structural perturbation information, thereby promoting more stable training and more efficient perturbation generation.
+Both parts are calculated using the adversarial loss defined in Eq.10.
+The intuition behind this design is that the gradient prior perturbation provides the generator with rich information about the prior gradient and perturbation structure, thereby promoting more stable training and more efficient perturbation generation.
+In particular, the optimized prior adversarial perturbation $\delta_{p}$ (noted as P-Pert) can be viewed as a cross-model universal adversarial perturbation that can protect multiple images from multiple facial manipulation models.
+The complete training process is given in Algorithm 1.
+input : Original dataset $S_{ori}$, facial manipulation model set $\mathcal{S}_{\mathcal{M}}$, max iteration $T$.
+output : Optimized $\mathcal{G}*$ and $\delta_{p}*$.
+1
+2// Generating initial prior adversarial perturbations $\delta_{p}^{0}$;
+3 $\delta_{p}^{0}\leftarrow 0$;
+4 $\hat{S}_{ori}\leftarrow SelectBatch(S_{ori})$;
+5 foreach $x^{i}\in\hat{S}_{ori}$ do
+6       // Implementing the PGD;
+7       $\delta^{i}\leftarrow PGD(x^{i})$;
+8       $\delta_{p}^{0}\leftarrow\delta_{p}^{0}+\delta^{i}$ ;
+9
+10 end foreach
+11
+[…]
+13
+[…]
+15 foreach $\lambda_{k}\in\mathcal{S}_{\lambda}^{0}$ do
+16       $\lambda_{k}^{0}\leftarrow 1$;
+17
+18 end foreach
+19// Training the Perturbation Generator;
+20 for $j\leftarrow 1$ to $T$ do
+21       $x_{b}\leftarrow SelectBatch(S_{ori})$;
+22       // Calculating the prior weight set;
+23       $\mathcal{S}_{\lambda}^{j}\leftarrow Strategy(\mathcal{G}^{j-1},\delta_{p}^{j-1},x_{b},\mathcal{L}^{adv},\mathcal{S}_{\mathcal{M}})$;
+24       // Updating the Parameters;
+25       $\mathcal{G}^{j},\delta_{p}^{j}\leftarrow Update(\mathcal{G}^{j-1},\delta_{p}^{j-1},x_{b},\mathcal{L}^{adv},\mathcal{S}_{\lambda}^{j},\mathcal{S}_{\mathcal{M}})$;
+26
+27 end for
+[…]
+29 $\delta_{p}*\leftarrow\delta_{p}^{N}$;
+Return $\mathcal{G}*,\delta_{p}*$
+Algorithm 1 Training of the Perturbation Generator
+III-E ID-Guard for Adversarial Training
+A potential application of ID-Guard is leveraging the proposed perturbation generator $\mathcal{G}$ as an adversarial attack module for adversarial training of facial manipulation models to enhance their robustness.
+During adversarial training, the well-trained $\mathcal{G}$ can rapidly generate training images with adversarial patterns, which are then used to train the facial manipulation model $\mathcal{M}$.
+Meanwhile, $\mathcal{G}$ is simultaneously fine-tuned to adapt to $\mathcal{M}$.
+Specifically, adversarial training can be formulated as a bi-level min-max optimization problem, where the inner maximization aims to generate adversarial examples that maximize the loss, while the outer minimization seeks to update the model parameters to minimize the bad-case loss.
+Given a facial manipulation model $\mathcal{M}$, it can be formulated as:
+$\min_{{\theta}_{\mathcal{M}}}\mathbb{E}_{(x,a)}\left[\max_{{\theta}_{\mathcal{G}},\mathcal{G}(x)\in\mathcal{S}}\mathcal{L}(\mathcal{M}(x+\mathcal{G}(x),a))\right],$ (15)
+where, the adversarial perturbation $\delta=\mathcal{G}(x)$ is constrained within a $\ell_{\infty}$-bounded perturbation set $\mathcal{S}$.
+The inner maximization problem optimizes the perturbation generator $\mathcal{G}$ to produce adversarial perturbations that maximize the loss.
+Specifically, the adversarial loss for the selected facial manipulation model $\mathcal{M}$ is computed in this step.
+The outer minimization then updates the parameters of $\mathcal{M}$ to enhance its robustness against such perturbations.
+Here, the training loss of $\mathcal{M}$ defined in the specific algorithm is computed.
+We will discuss its effectiveness in Section IV-D5.
 IV Experiments
@@
 IV-A1 Datasets
-In our experiments, the CelebAMask-HQ [47] dataset is selected for training the perturbation generator.
+In our experiments, the CelebAMask-HQ dataset [48] is selected to train the perturbation generator.
 It consists of more than 30,000 face images, where each image carries semantic masks for 19 facial component categories.
 These fine-grained mask labels can provide support for computing the binary mask loss during the training stage.
-To adequately evaluate the performance as well as the generalizability of our method and the competing algorithms, we test them on three datasets including CelebAMask-HQ [47], LFW [48] and FFHQ [49].
+To adequately evaluate the performance as well as the generalizability of our method and the competing algorithms, we test them on three datasets, including CelebAMask-HQ [48], LFW [49], and FFHQ [50].
+Figure 5: Visual example of $L_{2}^{face}$ metric design. (a) is a natural image, (b) is a forged image, and (c) and (d) are disrupted images in two different situations.
+Figure 6: Illustration of indicators that determine the success of a defense.
+(a) is a natural image, (b) is a forged image, (c), (d), and (e) are distorted images under different defense situations, respectively.
+It can be seen that when only one of the metrics, $L_{2}^{face}$ distance or identity similarity, satisfies the set conditions, it is not sufficient to break the identity of the individual in the image.
 IV-A2 Target Models
@@
 In the experiment, for StarGAN, AGGAN, FPGAN, and RelGAN, we select black hair, blond hair, brown hair, gender, and age as editing attributes; for HiSD, five images with black hair, blond hair, brown hair, glasses, and bangs are chosen as attribute references, respectively.
+TABLE I: Quantitative comparison for disrupting different target models.
+For each column within the same dataset.
+The best result is marked in bold, while the sub-optimal result is marked with an underline.
+Datasets Methods StarGAN[1] AGGAN[2] FPGAN[3] RelGAN[4] HiSD[5]
+[…]
+CelebAMask-HQ Disrupting[12] 1.047 0.023 1.000 0.114 0.479 0.292 0.134 0.369 0.472 0.021 0.753 0.001 0.004 0.839 0.001
+PG[10] 0.101 0.302 0.646 0.032 0.658 0.014 0.069 0.352 0.458 0.007 0.836 0.005 0.016 0.617 0.095
+CMUA[11] 0.586 0.368 0.584 0.062 0.646 0.016 0.052 0.486 0.126 0.296 0.630 0.070 0.055 0.603 0.165
+IAP[14] 0.450 0.118 0.994 0.054 0.300 0.398 0.321 0.181 0.928 0.109 0.545 0.165 0.056 0.193 0.590
+Ours (S-I) 0.362 0.055 1.000 0.376 0.062 1.000 0.558 0.004 1.000 0.285 0.065 0.998 0.205 0.018 1.000
+Ours (S-II) 0.592 0.016 1.000 0.298 0.088 1.000 0.635 0.001 1.000 0.404 0.013 1.000 0.204 0.043 0.998
+Ours (P-Pert) 0.246 0.031 1.000 0.288 0.078 1.000 0.551 0.009 1.000 0.069 0.176 0.770 0.136 0.171 0.902
+LFW Disrupting[12] 0.956 0.062 1.000 0.142 0.443 0.412 0.126 0.380 0.546 0.023 0.075 0.011 0.004 0.849 0.000
+PG[10] 0.134 0.306 0.728 0.053 0.656 0.044 0.069 0.415 0.356 0.008 0.838 0.002 0.020 0.651 0.035
+CMUA[11] 0.513 0.247 0.788 0.092 0.462 0.294 0.054 0.745 0.108 0.231 0.618 0.078 0.063 0.600 0.150
+IAP[14] 0.413 0.411 0.956 0.085 0.411 0.424 0.310 0.161 0.946 0.079 0.533 0.211 0.071 0.238 0.545
+Ours (S-I) 0.328 0.021 1.000 0.446 0.033 1.000 0.555 0.027 0.994 0.271 0.084 0.992 0.178 0.053 0.998
+Ours (S-II) 0.529 0.051 1.000 0.417 0.068 1.000 0.646 0.017 1.000 0.429 0.025 1.000 0.226 0.078 0.987
+Ours (P-Pert) 0.192 0.052 1.000 0.411 0.070 1.000 0.535 0.029 0.991 0.074 0.228 0.767 0.175 0.238 0.845
+FFHQ Disrupting[12] 0.956 0.033 1.000 0.142 0.487 0.312 0.126 0.409 0.426 0.023 0.747 0.013 0.005 0.878 0.002
+PG[10] 0.134 0.328 0.628 0.053 0.708 0.002 0.069 0.424 0.360 0.008 0.839 0.000 0.020 0.683 0.108
+CMUA[11] 0.515 0.346 0.624 0.092 0.635 0.028 0.054 0.512 0.168 0.231 0.666 0.043 0.063 0.656 0.120
+IAP[14] 0.413 0.167 0.976 0.085 0.365 0.476 0.310 0.207 0.898 0.079 0.568 0.123 0.071 0.252 0.525
+Ours (S-I) 0.328 0.053 1.000 0.446 0.089 0.994 0.557 0.019 1.000 0.271 0.112 0.973 0.178 0.050 0.965
+Ours (S-II) 0.534 0.005 1.000 0.330 0.103 0.970 0.509 0.023 1.000 0.348 0.028 0.995 0.184 0.082 0.963
+Ours (P-Pert) 0.213 0.052 1.000 0.301 0.090 0.978 0.455 0.046 1.000 0.063 0.246 0.637 0.118 0.262 0.730
 IV-A3 Baselines
-To demonstrate the superiority of the proposed method in face identity protection and cross-model universal performance, four advanced proactive defense methods including Disrupting [10], PG [39], CMUA [12] and IAP [15] are selected as competing algorithms.
+To demonstrate the superiority of the proposed method in face identity protection and cross-model transferable performance, four advanced proactive defense methods including Disrupting [12], PG [10], CMUA [11], and IAP [14], are selected as competing algorithms.
 Disrupting [12] disrupts facial manipulation by iteratively solving gradient-based adversarial perturbations on the target model.
-PG [39] achieves adversarial perturbation generation in gray-box scenarios by attacking a surrogate model.
+PG [10] achieves transferable adversarial perturbation generation in gray-box scenarios by attacking a surrogate model.
 CMUA [11] is a baseline of universal defense against multiple models.
-IAP [15] designs an information-containing adversarial perturbation, but we only implement its proactive distortion without involving the embedding and extraction of information to provide a fair comparison.
+IAP [14] designs an information-containing adversarial perturbation.
+For a fair comparison, we implement only its proactive disruption component, while the information embedding and extraction aspects are discussed in Section IV-E.
 IV-A4 Metrics
@@
 where $(i,j)$ is the coordinate of pixels and $Face_{i,j}$ is the binary facial mask of the image.
-The pixel value of its face area is 1, otherwise it is 0.
+The pixel value of its face area is $1$, otherwise, it is 00.
 The binary facial mask is calculated by Dlib  [^5] https://pypi.org/project/dlib.
+The intuition behind this design is that the traditional full-image $L_{2}$ distance metric fails to reflect distortions in the facial region.
+As illustrated in Fig.
+5, although the disrupted image in (c) is reported as successful under the $L_{2}$ distance metric, the perturbation primarily affects the background while leaving the facial region largely intact.
+In contrast, the proposed $L_{2}^{face}$ metric provides a more accurate measure of identity-related distortions in the facial region, which aligns with human perception.
 Additionally, we evaluate the identity similarity (noted as ID sim.
@@
 As shown in Fig.
-5, the distorted output in Fig.
-5 (c) reports a successful defense at the $L_{2}^{face}$ distance, but it appears that it merely “blackens” the face without destroying the individual’s identity.
+6, the distorted output in Fig.
+6 (c) reports a successful defense at the $L_{2}^{face}$ distance, but it seems to only blacken the face without destroying the individual’s identity.
 Therefore, we propose that both $L_{2}^{face}$ distance greater than $0.05$ and identity similarity less than $0.4$ be satisfied to indicate successful defense, which is a more challenging evaluation.
@@
 6 shows the necessity of considering both restrictions simultaneously.
-Figure 5: Illustration of indicators that determine the success of a defense.
-(a) is a natural image, (b) is a forged image, (c), (d), and (e) are distorted images under different defense situations, respectively.
-It can be seen that when only one of the metrics, $L_{2}^{face}$ distance or identity similarity, satisfies the set conditions, it is not sufficient to break the identity of the individual in the image.
 IV-A5 Implementation Details
@@
 The perturbation generator is trained using the Adam [52] with a learning rate of $0.0001$, and the batch size is $32$.
-TABLE I: Quantitative comparison for disrupting different target models.
-For each column within the same dataset.
-The best result is marked in bold, while the sub-optimal result is marked with an underline.
-  Datasets Methods StarGAN[21] AGGAN[22] FPGAN[23] RelGAN[24] HiSD[25]
-[…]
-CelebA-HQ Disrupting[10] 1.047 0.023 1.000 0.114 0.479 0.292 0.134 0.369 0.472 0.021 0.753 0.001 0.004 0.839 0.001
-PG[39] 0.101 0.302 0.646 0.032 0.658 0.014 0.069 0.352 0.458 0.007 0.836 0.005 0.016 0.617 0.095
-CMUA[12] 0.586 0.368 0.584 0.062 0.646 0.016 0.052 0.486 0.126 0.296 0.630 0.070 0.055 0.603 0.165
-IAP[4] 0.450 0.118 0.994 0.054 0.300 0.398 0.321 0.181 0.928 0.109 0.545 0.165 0.056 0.193 0.590
-Ours (S-I) 0.362 0.055 1.000 0.376 0.062 1.000 0.558 0.004 1.000 0.285 0.065 0.998 0.205 0.018 1.000
-Ours (S-II) 0.587 0.018 1.000 0.300 0.089 1.000 0.631 0.000 1.000 0.408 0.010 1.000 0.202 0.042 0.998
-  LFW Disrupting[10] 0.956 0.062 1.000 0.142 0.443 0.412 0.126 0.380 0.546 0.023 0.075 0.011 0.004 0.849 0.000
-PG[39] 0.134 0.306 0.728 0.053 0.656 0.044 0.069 0.415 0.356 0.008 0.838 0.002 0.020 0.651 0.035
-CMUA[12] 0.513 0.247 0.788 0.092 0.462 0.294 0.054 0.745 0.108 0.231 0.618 0.078 0.063 0.600 0.150
-IAP[4] 0.413 0.411 0.956 0.085 0.411 0.424 0.310 0.161 0.946 0.079 0.533 0.211 0.071 0.238 0.545
-Ours (S-I) 0.328 0.021 1.000 0.446 0.033 1.000 0.555 0.027 0.994 0.271 0.084 0.992 0.178 0.053 0.998
-Ours (S-II) 0.522 0.047 1.000 0.414 0.078 1.000 0.644 0.017 1.000 0.430 0.022 1.000 0.221 0.081 0.980
-  FFHQ Disrupting[10] 0.956 0.033 1.000 0.142 0.487 0.312 0.126 0.409 0.426 0.023 0.747 0.013 0.005 0.878 0.002
-PG[39] 0.134 0.328 0.628 0.053 0.708 0.002 0.069 0.424 0.360 0.008 0.839 0.000 0.020 0.683 0.108
-CMUA[12] 0.515 0.346 0.624 0.092 0.635 0.028 0.054 0.512 0.168 0.231 0.666 0.043 0.063 0.656 0.120
-IAP[4] 0.413 0.167 0.976 0.085 0.365 0.476 0.310 0.207 0.898 0.079 0.568 0.123 0.071 0.252 0.525
-Ours (S-I) 0.328 0.053 1.000 0.446 0.089 0.994 0.557 0.019 1.000 0.271 0.112 0.973 0.178 0.050 0.965
-Ours (S-II) 0.536 0.002 1.000 0.329 0.107 0.970 0.502 0.025 1.000 0.345 0.030 0.995 0.183 0.081 0.964
+Figure 7: Visual examples of disruption to different facial manipulations.
 IV-B Comparison with Baselines
@@
 Our method is reported separately under two strategies, as presented in III-C.
-The Disrupting [10] produces gradient-based adversarial perturbations against StarGAN [21], which makes it overfit in disrupting this model at the cost of cross-model performance.
+The Disrupting [12] produces gradient-based adversarial perturbations against StarGAN [1], which makes it overfit in disrupting this model at the cost of cross-model performance, as shown in Fig.
+7.
 The perturbation optimization of CMUA [11] and PG [10] is unconstrained and thus has limited destruction to the identity semantics.
-As shown in Fig.
-6, the person’s identity in a distorted image under the CMUA can be still recognized.
-Attributed to the feature correlation measurement loss employed, IAP [15] improves the performance to destroy identification to some limit extent.
+The visual example of the CMUA in Fig.
+7 shows that the person’s identity in the distorted image can still be recognized.
+Attributed to the feature correlation measurement loss employed, IAP [14] improves the performance to destroy identification to some extent.
 In comparison, our method significantly destroys the identity semantics of images and effectively prevents the stigmatization of faces.
@@
 Due to the introduction of the dynamic parameter strategy, we achieve balanced performance against various facial manipulations.
-For the most robust model against attack, HiSD [25], ID-guard improves the defense success rate by 50% compared with the state-of-the-art method.
+For the most robust model against attack, HiSD [5], ID-Guard improves the defense success rate by $50\%$ compared with the state-of-the-art method.
 On the three selected datasets, ID-Guard under both strategies demonstrated superior performance.
-For strategy I, the MGDA algorithm is employed to adjust the weights, which has the advantage of completely eliminating the need for human intervention and prior knowledge during the training process.
-However, this non-intervention makes ID-guard based on strategy I exhibit an extreme effect to some extent: it performs better on the more vulnerable model (e.g., AGGAN [22]) and the more robust model (e.g., HiSD [25]), but does not provide significant improvement on models in the middle (e.g., RelGAN [24]).
+For strategy I, the MGDA algorithm is employed to adjust the weights, which has the advantage of eliminating the need for human intervention and prior knowledge during the training process.
+However, this non-intervention makes MGDA-based ID-Guard exhibit an extreme effect to some extent: it performs better on the more vulnerable model (e.g., AGGAN [2]) and the more robust model (e.g., HiSD [5]), but does not provide significant improvement on models in the middle (e.g., RelGAN [4]).
 Moreover, additional backpropagation computation is required at each iteration when implementing MGDA, which increases the training overhead.
-In contrast, Strategy II-based ID-guard only additionally computes a set of KPI values, which has a minimal impact on training overhead.
-Strategy II maintains a set of prior parameters for balancing each model, thereby further balancing performance.
-The disadvantage of it is the extra work of determining the prior parameter set.
-Figure 6: Visual examples of disruption to different facial manipulations.
-Compared with CMUA [12], which still retains a large amount of identity semantics in the distorted image, our method makes the face in the image completely unrecognizable.
+In contrast, the KPI-based strategy balances different models by maintaining a set of prior parameters, ensuring more stable defense performance across models and a more balanced performance improvement.
+However, its drawback is that preliminary experiments are required before formal training to determine the values of the prior weight set.
+We conducted an additional evaluation of the optimized gradient-prior perturbation, P-Pert.
+Similar to CMUA, P-Pert is a cross-model universal adversarial perturbation designed to protect multiple facial images.
+As shown in the quantitative results Table I and the visualizations in Fig.
+7, P-Pert significantly outperforms the baseline universal perturbation in both facial region disruption and cross-model balance.
+This advantage stems from its use of a dynamic weight strategy and an identity disruption module for optimization.
+Compared to the proposed generative perturbation, P-Pert achieves cross-image universality at the cost of a certain degree of reduced defense performance.
+Defenders can balance performance and computational efficiency by selecting either $\mathcal{G}$ or P-Pert for image protection.
 IV-C Ablation Study
@@
 As shown in Fig.
-7, the mask loss concentrates the distortion on the face region of the image, while the identity loss destroys the global texture.
+8, the mask loss concentrates the distortion on the face region of the image, while the identity loss destroys the global texture.
 Feature loss brings overall gain, which benefits from the similarity in feature extraction of the face manipulation model.
@@
 #1 w/o all 0.172 0.509 0.358
-#2 w/o Feature loss 0.396 0.050 0.973
-#3 w/o Id. loss 0.431 0.148 0.937
-#4 w/o Mask loss 0.189 0.039 0.866
+#2 w/o FCL 0.396 0.050 0.973
+#3 w/o ICL 0.431 0.148 0.937
+#4 w/o MCL 0.189 0.039 0.866
 #5 w/ all 0.425 0.031 0.999
+          FCL means the Feature Confusion Loss.
+          ICL means the Identity Consistency Loss.
+          MCL means the Mask Constrained Loss.
 Figure 8: Visual examples of ablation study of Identity Destruction Module.
@@
 Although HMM balances each model to a certain extent, it ignores the difference in model gradients and thus causes the degradation of average performance.
-Separate prior weight setting or KPI are unstable and difficult to set, so we cleverly blend the two in Strategy II and get stable training.
+Separate prior weight settings or KPI are unstable and difficult to set, so we cleverly blend the two in Strategy II and get stable training.
 The benefits of this are two-fold: 1) It reduces the difficulty of a prior setting, and only needs to determine a series of orders of magnitude to allow automatic optimization of parameters; 2) It makes the KPI strategy more stable.
@@
 DTP 1.000 0.894 1.000 0.802 0.885 0.916
-Ours (S-I) 1.000 1.000 1.000 0.998 1.000 0.999
-Ours (S-II) 1.000 1.000 1.000 1.000 0.998 0.999
+MGDA-based (S-I) 1.000 1.000 1.000 0.998 1.000 0.999
+KPI-based (S-II) 1.000 1.000 1.000 1.000 0.998 0.999
 IV-C3 Gradient Prior Perturbation
 Gradient prior perturbation aims to provide the generator with noise-like prior knowledge, thus accelerating its convergence.
-For comparison, the variation of training loss and defense performance at different scales with gradient a prior perturbation, with a prior random noise, and without prior knowledge are shown in Fig.
-8.
+For comparison, the variation of training loss and defense performance at different scales with the gradient prior perturbation, with a prior random noise, and without prior knowledge is shown in Fig.
+9.
 Both gradient prior perturbation and random noise promote the convergence of the generator, which is due to the introduction of global noise structure [9].
-In terms of generator performance, methods based on gradient prior perturbation at different scales have shown the most significant defense effect, with an average improvement of 31.2% compared to random noise methods.
-We believe that the reason behind this is that gradient prior perturbation involves rich adversarial structural information.
+In terms of generator performance, methods based on gradient prior perturbation at different scales have shown the most significant defense effect, with an average improvement of $31.2\%$ compared to random noise methods.
+We believe that the reason behind this is that the gradient prior perturbation involves rich adversarial structural information.
 (a) Changes of training loss
@@
 We explore the impact of different generator architectures on performance.
-Three mainstream architectures including Unet [52], Resnet [38] and Transformer [53] were selected as the generators of the proposed ID-guard.
+Three mainstream architectures, including Unet [43], Resnet [51], and Transformer [55], are selected as the generators of the proposed ID-Guard.
 Fig.
@@
 Compared with Unet, Resnet and Transformer architectures have achieved significant advantages.
-As shown in Table IV, Transformer achieved optimal performance at the expense of model parameter size, while Resnet achieved very close performance with less than 5% of its parameter size.
+As shown in Table IV, Transformer achieved optimal performance at the expense of model parameter size, while Resnet achieved very close performance with less than $5\%$ of its parameter size.
 We propose to use Resnet as the architecture for the generator of the proposed ID-Guard, and the intuition behind this is that the generated perturbation can be regarded as a residual of the image.
@@
 Therefore, the threat of stigmatization of distorted images comes not only from the human eye but also from commercial facial recognition systems.
-Therefore, the threat of stigmatization of distorted images comes not only from the human eye but also from commercial facial recognition systems.
 As shown in Fig.
-10, we evaluate the misdirection success rates of the destroyed outputs of ID-guard and competing algorithms on three mainstream face recognition systems.
-As can be seen, our method reports optimal results, achieving over 95% misdirection success rate on Google  [^6] https://www.google.com and StarByFace  [^7] https://starbyface.com.
-Baidu  [^8] https://www.baidu.com has the most robust recognition system, with CMUA [12] and PG [39] can hardly fool it, but ID-guard still causes it to recognize more than 75% of images incorrectly.
+11, we evaluate the misdirection success rates of the destroyed outputs of ID-Guard and competing algorithms on three mainstream face recognition systems.
+As can be seen, our method reports optimal results, achieving over $95\%$ misdirection success rate on Google  [^6] https://www.google.com and StarByFace  [^7] https://starbyface.com.
+Baidu  [^8] https://www.baidu.com has the most robust recognition system, with CMUA [11] and PG [10] can hardly fool it, but ID-Guard still causes it to recognize more than $75\%$ of images incorrectly.
 The good performance of ID-Guard is due to the identity loss introduced to destroy the identity recognition baseline model, which is widely used in commercial face recognition systems.
 Figure 11: Quantitative comparison of the misdirection success rates of distorted images on three mainstream commercial face recognition systems.
-IV-D2 Resisting Face Inpaintings
+IV-D2 Resisting facial inpainting
 Another challenge comes from the image inpainting system.
-A well-trained face inpainting model can recover distorted facial images, rendering defenses ineffective.
-We evaluate the performance of distorted images against two baseline face inpainting models, namely LBP [54] and GS-SSA [55].
+A well-trained facial inpainting model can recover distorted facial images, rendering defenses ineffective.
+We evaluate the performance of distorted images against two baseline facial inpainting models, namely LBP [53] and GS-SSA [54].
 The quantitative results of $L_{2}^{face}$ distance are reported in Table V, and the visualization results are shown in Fig.
-11.
+12.
 Although the difference between the repaired distorted image and the forged result is greatly reduced, the proposed method still exhibits optimal defense performance.
-This is because these face inpainting systems rely heavily on undistorted regions of the image, However, we achieve a greater degree of destruction of the entire image texture due to the introduction of feature loss.
+This is because these facial inpainting systems rely heavily on undistorted regions of the image, However, we achieve a greater degree of destruction of the entire image texture due to the introduction of feature loss.
 TABLE V: Quantitative results of the image inpainting on distorted images. The best result is marked in bold, while the sub-optimal result is marked with an underline.
-  Face inpaintings Methods StarGAN AGGAN FPGAN RelGAN HiSD
+facial inpainting Methods StarGAN AGGAN FPGAN RelGAN HiSD
 LBP [53] Disrupting[12] 0.664 0.114 0.156 0.063 0.053
@@
 Ours (S-I) 0.259 0.141 0.335 0.197 0.084
-Ours (S-II) 0.387 0.140 0.340 0.214 0.075
+Ours (S-II) 0.389 0.137 0.344 0.218 0.076
 GS-SSA [54] Disrupting[12] 0.772 0.081 0.106 0.028 0.017
@@
 Ours (S-I) 0.296 0.108 0.328 0.143 0.059
-Ours (S-II) 0.478 0.110 0.374 0.149 0.056
+Ours (S-II) 0.481 0.113 0.372 0.150 0.058
 Figure 12: Visual examples of the image inpainting on distorted images. The original and fake images are provided on the right side for reference.
@@
 Gray box scenarios are defined where the model type is known but the internal parameters are not accessible.
-We train a StarGAN [21] and RelGAN [24] as the target gray-box model respectively.
-As shown in Table VI, to some extent, ID-guard still maintains the most powerful defense capabilities.
-Furthermore, it can be found that compared to competing algorithms, ID-guard has the smallest performance degradation.
+We train a StarGAN [1] and RelGAN [4] as the target gray-box model, respectively.
+As shown in Table VI, Disrupting derives adversarial perturbations on the accessible white-box versions of StarGAN and RelGAN, respectively.
+This model-specific perturbation generation enables it to maintain gray-box transferability to a certain extent.
+ID-Guard still maintains the most powerful defense capabilities.
 We attribute this to the designed feature loss, which destroys model feature extraction with cross-model consistency [25].
+Additionally, the baseline methods perform better against the gray-box RelGAN than against its white-box version.
+The reason for this could be that their imbalanced training process caused the perturbation performance to be biased toward the most vulnerable white-box StarGAN, which has adversarial gradients that are closer to the gray-box RelGAN used in this experiment.
 TABLE VI: Quantitative results in gray box scenarios. The best result is marked in bold, while the sub-optimal result is marked with an underline.
@@
 CMUA[11] 0.090 0.798 0.035 0.041 0.719 0.008
-IAP[4] 0.106 0.560 0.122 0.114 0.291 0.653
-Ours (S-I) 0.129 0.363 0.584 0.107 0.294 0.693
-Ours (S-II) 0.109 0.351 0.532 0.124 0.111 0.890
+IAP[14] 0.088 0.573 0.117 0.096 0.372 0.421
+Ours (S-I) 0.137 0.305 0.620 0.168 0.207 0.724
+Ours (S-II) 0.133 0.329 0.581 0.176 0.100 0.906
 indicates the facial manipulation model in the gray-box setting.
-IV-D4 Robustness in Real Social Scenarios
+IV-D4 Robustness under Lossy Operations
 In real scenarios, users often upload perturbed images to social applications to share their lives.
 However, various lossy operations on the transmission channel can destroy the effectiveness of the perturbation.
-In this section, we first evaluate the robustness of competing algorithms on Facebook and WeChat.
-Next, we incorporate the robustness strategy in [18] into our generator training, which can be viewed as a downstream task of this work.
-Quantitative results are shown in Table VII.
-Even without robustness training, our method still reports the best robustness compared to the baseline.
-The reason may be that our adversarial perturbation is not averaged over the image, but is constrained to focus on faces, which makes it harder to neutralize.
+In this section, we evaluate the robustness of ID-Guard and competing algorithms under JPEG compression and Gaussian blur.
+To further verify the integration capability of our framework and the robustness strategy, for JPEG, we incorporate the compression-resistant strategy from [27] into our generator training.
+The results are reported as “Ours (S-II) w/ R.”.
 As shown in Fig.
-12, when the robust training strategy is integrated, the anti-lossy operation performance of generated perturbations is greatly improved.
-This demonstrates the flexibility of the proposed ID-guard framework to effectively integrate with progressive strategies in the community.
-TABLE VII: Quantitative results on the robustness of adversarial perturbations in real scenarios.
-The adversarial images are uploaded to designated social applications and then downloaded.
-The best result is marked in bold.
-“w/” and “w/o” indicate whether robust training is used or not, respectively.
-[…]
-Facebook Disrupting[10] 0.023 0.760 0.022
-PG[39] 0.012 0.744 0.009
-CMUA[12] 0.003 0.761 0.001
-IAP[4] 0.012 0.749 0.016
-Ours (S-II) w/o 0.041 0.708 0.112
-Ours (S-II) w/ 0.065 0.677 0.306
-  WeChat Disrupting[10] 0.010 0.762 0.006
-PG[39] 0.004 0.747 0.003
-CMUA[12] 0.002 0.741 0.001
-IAP[4] 0.004 0.746 0.001
-Ours (S-II) w/o 0.017 0.739 0.024
-Ours (S-II) w/ 0.052 0.702 0.134
-Figure 12: Visual examples of defense effects in real scenarios.
-The transmission channel is Facebook and the StarGAN is chosen as the target model.
-“w/” and “w/o” indicate whether robust training is used or not, respectively.
+13, as the intensity of the lossy operation increases, the defense performance of each method is gradually weakened.
+Our method demonstrates significant robustness at different scales.
+The underlying reason for this could be the introduction of the identity disruption module, which concentrates the effectiveness of the adversarial perturbations in specific areas, making them less susceptible to degradation from lossy operations.
+When the robustness training strategy was integrated, ID-Guard’s robustness improved significantly.
+A sample visualization is shown in Fig.
+14.
+This demonstrates the flexibility of the proposed framework and its ability to effectively integrate with advanced strategies from the research community.
+TABLE VII: Quantitative results of the adversarial robustness of the original face manipulation model and its adversarial training version against different attacks.
+Model Gaussian noise ($\sigma=0.05$) FGSM I-FGSM PGD C&W
+$L_{2}^{face}$downarrow ID sim.$\uparrow$ $L_{2}^{face}$$\downarrow$ ID sim.$\uparrow$ $L_{2}^{face}$$\downarrow$ ID sim.$\uparrow$ $L_{2}^{face}$$\downarrow$ ID sim.$\uparrow$ $L_{2}^{face}$$\downarrow$ ID sim.$\uparrow$
+StarGAN 0.070 0.261 0.185 0.184 1.075 0.069 1.102 0.004 1.377 0.009
+StarGAN-AT 0.001 0.985 0.011 0.977 0.148 0.700 0.141 0.712 0.149 0.756
+RelGAN 0.003 0.905 0.140 0.766 0.584 0.186 0.615 0.145 0.983 0.099
+RelGAN-AT 0.000 0.992 0.001 0.989 0.019 0.783 0.018 0.791 0.002 0.988
+(a) JPEG compression
+(b) Gaussian Blur
+Figure 13: The performance of the defense algorithm against StarGAN under different intensities of lossy operations.
+In the JPEG evaluation, the performance of ID-Guard integrated with the compression-resistant strategy has a significant advantage.
+Therefore, in order to present the results, we take the logarithm of 10 for $L_{2}^{face}$.
+Figure 14: Visualization of adversarial examples against the StarGAN model under JPEG compression with QF=85.
+IV-D5 ID-Guard for Adversarial Training
+In this section, we evaluate the effectiveness of using the ID-Guard framework for adversarial training of facial manipulation models.
+Specifically, we select StarGAN[1] and RelGAN[4] as example facial manipulation models and apply the adversarial training (AT) approach proposed in Section III-E to obtain StarGAN-AT and RelGAN-AT.
+The adversarial training version of the facial manipulation model $\mathcal{M}$ is noted as $\mathcal{M}^{AT}$.
+The training framework follows their official open-source implementations  [^9] https://github.com/yunjey/stargan, https://github.com/elvisyjlin/RelGAN-PyTorch.
+To evaluate adversarial robustness, we test these models against Gaussian noise ($\sigma=0.05$), FGSM[36], I-FGSM[66], PGD[37], and C&W[65] attacks.
+All attacks follow the white-box setup.
+As reported in Table VII, non-adversarial Gaussian noise and the single-step FGSM disrupt $\mathcal{M}$ to some extent but have almost no impact on $\mathcal{M}^{AT}$.
+Furthermore, for the three stronger attack algorithms, I-FGSM, PGD, and C&W, the model implemented adversarial training demonstrates significant robustness.
+Especially for RelGAN-AT, the $L_{2}^{face}$ consistently remains within the threshold of $0.05$.
+Additionally, the quantitative identity similarity results and the visualized examples in Fig.
+15 indicate that these adversarial attacks mainly distort the background regions of $\mathcal{M}$’s manipulated outputs while leaving the facial regions unaffected.
+This validates that the proposed ID-Guard framework can serve as a plug-and-play adversarial attack module within adversarial training, significantly enhancing the robustness of facial manipulation models.
+Figure 15: Visual example of the adversarial robustness of the original face manipulation model and its adversarial training version against PGD[37].
 IV-E Further Discussion
@@
 Here, to explore its mechanism in depth, we record the dynamic changes of the weight set $\mathcal{S}_{\lambda}=\left\{\lambda_{1},\lambda_{2},\dots,\lambda_{N}\right\}$ of attack losses in an epoch training, as shown in Fig.
-13.
-For strategy I, we use the MGDA algorithm to solve the set of weights automatically in each iteration.
+16.
+For strategy I, we leverage the MGDA algorithm to optimize the set of weights automatically in each iteration.
 It can be found that face manipulation models with strong robustness, such as HiSD [5] and RelGAN [4], tend to be assigned larger weights to strengthen the attack against them.
@@
 Figure 16: The changing trend of the weight of attack loss with the number of iterations for different facial manipulations when adopting the dynamic weight strategy.
-IV-E2 What does the generator learn?
+IV-E2 Can ID-guard be migrated to active forensics?
+We migrate the proposed ID-Guard framework to proactive forensics for facial manipulation to demonstrate its scalability.
+Specifically, we follow the setup in [14]: Given a natural image $x$ and a binary message $m\in\{0,1\}^{L}$ of length $L$, a message embedding encoder $E_{m}$ maps the image and a corresponding message to a watermarked image.
+After the watermarked image undergoes manipulation, a message extractor decoder $D_{m}$ is used to recover $m$ from the forged image.
+The training objective is to minimize $\sum_{k=1}^{N}L^{wm}_{k}=MSE(m,D_{m}(\mathcal{M}_{k}(E_{m}(x,m))))$, while an additional discriminator is employed to ensure the visual quality of the watermarked image.
+FaceTagger[59] and information-based IAP[14] are selected as baselines.
+In our method, the proposed dynamic weight strategy is applied to weight $L^{wm}_{k}$ for different $\mathcal{M}_{k}$ during training.
+The Bit Error Rate (BER) and the visual quality of the watermarked image are quantitatively reported in Table VIII.
+FaceTagger is trained with StarGAN[1] as the target model, leading to overfitting and consequently poor overall performance.
+Compared to IAP, our method demonstrates more balanced forensic performance across different facial manipulation models, particularly against HiSD[5], which exhibits significant structural differences.
+This result highlights the strong scalability and transferability of the proposed framework, making it a plug-and-play tool adaptable to various cross-model tasks in the community.
+TABLE VIII: Quantitative results of proactive forensics. The best result is marked in bold.
+[…]
+StarGAN AGGAN FPGAN RelGAN HiSD
+FaceTagger[59] 0.000 0.501 0.009 0.069 0.512 0.995
+IAP[14] 0.034 0.144 0.000 0.001 0.000 0.983
+Ours 0.000 0.000 0.000 0.000 0.000 0.995
+IV-E3 What does the generator learn?
 The adversarial perturbations generated by the generator trained under different loss constraints are shown in Fig.
-14.
-The mask loss concentrates rich adversarial information on the facial area of the image, thereby completely distorting the output face.
+17.
+The mask loss concentrates rich adversarial information on the facial area of the image, thus completely distorting the output face.
 The generator trained using only the identity loss learns to disrupt images at the texture level.
-Combined with the results in Fig 7, this destruction changes the key feature semantics and visual attributes of the face.
+Combined with the results in Fig.
+8, this destruction changes the key feature semantics and visual attributes of the face.
 Therefore, the standard generator learns to generate adversarial perturbations that cause maximum damage to facial regions and alter the identifiable texture features of the image.
 Figure 17: Visual examples of the generated perturbation.
-Among them, (a) is the original image, (b) is the adversarial image, (c) is the perturbation generated by the generator trained only with mask loss, (d) is the perturbation generated by the generator trained only with identity loss, and (e) is the perturbation generated by the standard generator.
+Among them, (a) is the original image, (b) is the adversarial image, (c) is the perturbation generated by the generator trained only with mask constrained loss, (d) is the perturbation generated by the generator trained only with identity consistency loss, and (e) is the perturbation generated by the standard generator.
 V Conclusion
-In this work, we proposed a universal adversarial framework for combating facial manipulation, named ID-guard.
+In this work, we proposed a universal framework for combating facial manipulation, named ID-Guard.
 To prevent face stigmatization problems caused by unconstrained image distortion, we propose an Identity Destruction Module to eliminate identity semantics.
 Furthermore, to improve the cross-model performance of generating perturbations, we regard attacking different models as a multi-task learning problem and introduce a dynamic parameter strategy.
-The proposed method not only effectively resists multiple face manipulations, but also significantly disrupts face identification.
+The proposed method not only effectively resists multiple face manipulations but also significantly disrupts face identification.
 In addition, the experiment also demonstrated the possibility of ID-Guard in circumventing commercial face recognition systems and image inpaintings.